### v1

# Graph Neural Network Training Systems: A Performance Comparison of Full-Graph and Mini-Batch. [Experiment, Analysis & Benchmark]

Abstract.
Graph Neural Networks (GNNs) have gained significant attention in recent years due to their ability to learn representations of graph-structured data. Two common methods for training GNNs are mini-batch training and full-graph training. Since these two methods require different training pipelines and systems optimizations, two separate classes of GNN training systems emerged, each tailored for one method. Works that introduce systems belonging to a particular category predominantly compare them with other systems within the same category, offering limited or no comparison with systems from the other category. Some prior work also justifies its focus on one specific training method by arguing that it achieves higher accuracy than the alternative. The literature, however, has incomplete and contradictory evidence in this regard.
In this paper, we provide a comprehensive empirical comparison of representative full-graph and mini-batch GNN training systems. We find that the mini-batch training systems consistently converge faster than the full-graph training ones across multiple datasets, GNN models, and system configurations, with speedups between 2.4×2.4$\times$2.4 × - 15.2×15.2$\times$15.2 ×. We also find that both training techniques converge to similar accuracy values. Our work highlights the importance of comparing system across different classes, using time-to-accuracy rather than epoch time for performance comparison, and selecting appropriate hyperparameters for the comparison.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/saurabhbajaj123/GNN_minibatch_vs_fullbatch.
1. Introduction
Graph neural networks (GNNs) are a class of machine learning models that reached state-of-the-art performance in many tasks related to the analysis of graph-structured data, including social network analysis, recommendations, and fraud detection (Zitnik et al., 2018; Schlichtkrull et al., 2018; Hamilton et al., 2018). They are often used to process large graphs that have millions of vertices and billions of edges (Ying et al., 2018; Hu et al., 2021; Liu et al., 2023). A large volume of recent work, both in academia and industry, has been dedicated to scaling GNN training to support such large graphs using multi-GPU systems. This is a challenging problem because GNNs run multiple rounds of message passing across neighboring vertices, which is an irregular computation.
Two classes of GNN systems: Full-graph and mini-batch. GNNs can be trained using either a mini-batch or a full-batch (typically called full-graph) approach, much like other machine learning models. In GNN training, however, these two approaches result in different message-passing patterns, which require different system pipelines and pose different system optimization challenges. This resulted in the development of two distinct classes of GNN training systems, each designed to support scalability for either mini-batch or full-graph training.
Figure 1. Different data management pipelines in two example systems: PipeGCN (full-graph) and DGL (mini-batch). The diagrams omit gradient synchronization.
The main reason for the emergence of these two separate class of systems is that different training methods (full-graph vs. mini-batch) require different data management pipelines to partition data, transfer it across CPUs and GPUs, and parallelize computation when scaling to large graphs. Full-graph GNN training performs message-passing for the entire batch, typically across the entire graph, at each epoch. To scale to large graphs that do not fit in the memory of a single GPU, multi-GPU full-graph training systems use model parallelism: they partition the graph, process different partitions at different GPUs, and perform message passing across partitions (Tripathy et al., 2020; Ma et al., 2019; Jia et al., 2020). For example, the PipeGCN (Wan et al., 2022b) full-graph training system partitions the input graph and keeps each partition in a different GPU (see Figure 1a). The black dotted lines show the communication of vertex features across GPUs during message passing at each GNN layer in the forward and backward pass. The communication happens through either a fast cross-GPU bus such as NVLink, if available, or PCIe.
Mini-batch training breaks up the training set into mini-batches. Each training epoch now consists of multiple iterations, one per mini-batch. Mini-batch training is popular in production systems because it is amenable to scaling through data parallelism (Zheng et al., 2020, 2022b; Zhu et al., 2019). A unique challenge when applying data parallelism to GNNs is that preparing mini-batches requires expensive pre-processing at each iteration. At each iteration, data-parallel mini-batch training systems sample the k-hop neighborhood of the vertices in the mini-batch, load each sample into a different GPU and perform local message passing only within each sample, without passing message across GPUs. For example, Figure  1b shows a training iteration in the popular DGL system. First, each GPU obtains a sample of the k-hop neighborhood of a subset of the vertices in the mini-batch. Sampling is preferably executed using the GPU since it is a computationally expensive step in itself (Jangda et al., 2021). Next, the sample and the input features of its vertices are loaded to the GPUs. Finally, each GPU performs a forward and backward pass on the local sample independently.
These different data management pipelines result in a different set of optimizations. Full-graph training systems must deal with expensive and irregular computation and high communication costs at each epoch. They propose sophisticated scheduling and partitioning algorithms to address these bottlenecks (Md et al., 2021; Peng et al., 2022; Wan et al., 2022b; Thorpe et al., 2021; Wan et al., 2023a). Work can be partitioned and scheduled offline, since the graph is static. Some full-graph training systems also propose techniques to reduce the volume of message data transferred across GPUs and hosts (Wan et al., 2023b; Peng et al., 2022; Wan et al., 2022a). In data-parallel mini-batch training systems, instead, computation at each iteration is much more lightweight and there is no cross-GPU message passing. However, pre-processing is a major bottleneck. Samples are typically large, so sampling them and loading them into GPUs is expensive. Research on scalable mini-batch systems has focused on addressing these bottlenecks by accelerating mini-batch sampling (Yang et al., 2022b; Jangda et al., 2021; Pandey et al., 2020; Wang et al., 2021a; Gong et al., 2023; Cai et al., 2023; Zheng et al., 2022a, b) and reducing the cost of loading the input features by caching them in the memory of the GPUs (Lin et al., 2020; Yang et al., 2022b; Song et al., 2023b; Cai et al., 2023; Tan et al., 2023; Yang et al., 2022a; Liu et al., 2023; Polisetty et al., 2023).
Motivation: What is the state of the art? To understand the state of the art in the field it is necessary to compare systems across different classes. We performed an exhaustive review of the existing literature and found that prior works that propose systems in one class compare them with other systems in the same class, with limited or no performance evaluation of systems in the other class. Furthermore, several papers assert the superiority of the training approach they support in terms of the best accuracy that can be achieved, resulting in conflicting claims. We report a detailed discussion of the evidence in the literature in Section  2. Overall, the existing literature makes it difficult to get a clear picture of the state of the art of the field.
Performance evaluation. In this paper, we address this open issue by performing a thorough performance comparison of representative state-of-the-art GNN training systems across the two classes (Wan et al., 2022b, a; Wang et al., 2019; Zheng et al., 2020; Tan et al., 2023). We chose systems that have a stable, publicly-available, and open-source implementation, are representative of the most important types of systems optimizations in their class, and implement all the standard GNN models (GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GAT (Veličković et al., 2018)) in their repository.
Our results show that the mini-batch training systems are consistently faster than the full-graph training counterparts in reaching the same target accuracy across all datasets, GNN model, and hardware deployments we considered. These results are consistent with the evidence in the literature. Several papers that propose full-graph training systems use epoch time as the main metric to compare with mini-batch training systems, showing that the latter have a larger epoch time  (Wang et al., 2022; Wan et al., 2023a; Md et al., 2021; Chen et al., 2023). Our evaluation confirms these results. However, our results also highlight that considering the epoch time alone only gives a partial picture and that time-to-accuracy is a better metric for comparison. Mini-batch training typically requires fewer epochs to converge because it updates the model parameters multiple times per epoch, once at each iteration, whereas full-graph training performs only one update per epoch. Therefore, even if mini-batch training systems perform more work per epoch than full-graph training ones to run multiple iterations, their overall time-to-accuracy is lower. Our empirical results show that this holds consistently across all systems, deployments, models, and datasets we consider. Across all the datasets we considered, the fastest mini-batch training system is between 2.4×2.4$\times$2.4 × to 15.2×15.2$\times$15.2 × (6.3×6.3$\times$6.3 × on average) faster than the fastest full-graph systems in terms of time-to-accuracy.
In terms of scalability, full-graph training and mini-batch training systems observe similar relative speedups when they scale from a single GPU to multiple GPUs. However, the mini-batch systems we consider have a faster time-to-accuracy on a single GPU or host, so they remain faster in absolute terms in all configurations. We report our performance results in Section 4.
Performance vs. accuracy. The time-to-accuracy experiments discussed previously measure the time it takes different systems to reach the same target accuracy, which we select by performing hyperparameter tuning, using the same model and dataset. However, using a system with a longer time-to-accuracy can be justifiable if a model can converge to a higher accuracy. This is an important factor when navigating the performance trade-offs of using different GNN training systems. In fact, some papers proposing full-graph GNN systems assert that full-graph training achieves higher accuracy (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others proposing mini-batch training systems claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
Our results show that, with proper hyperparameter tuning, both full-graph and mini-batch training reach a similar accuracy across different datasets and models. Which method achieves the highest accuracy depends on the dataset and GNN model but the accuracy gaps are small. In our evaluation, we matched or exceeded all the accuracy results in the literature on GNN training systems we reviewed. We observe that given a dataset and GNN model, if we use the best hyperparameters for one training method with a system using the other training method, the latter achieves lower accuracy. This dependence on the choice of hyperparameters may explain the contradicting results found in the literature. Our results underline the importance of conducting separate hyperparameter tuning for the two training methods when comparing performance and accuracy across systems using different training methods. We also find that finding good hyperaparameters is easier with mini-batch training because it is less sensitive to variations in the model architecture. This has an impact on the overall training time, since model training often involves performing hyperparameter search. The harder it is to find good hyperparameters, the longer the overall training time. We report our accuracy experiments in Section 5.
Lessons learned. Our results show that the mini-batch training systems we consider consistently achieve a better performance than the full-graph training ones and a similar accuracy. We stress, however, that the goal of this study is not to establish the inherent superiority of one class of systems over the other. Our study is an empirical evaluation and the field is continuously evolving.
From a methodological perspective, our work highlights the importance of comparing GNN training systems across different classes. It also indicates a set of principles on how to perform this comparison: use time-to-accuracy as primary metric for comparison rather than epoch time, and perform separate hyperparameter tuning for mini-batch and full-graph training when selecting a target accuracy for performance comparison or comparing convergence accuracy. Prior works do not always follow these principles, and this may explain the apparent contradiction in the existing results. Our work offers guidance to ML practitioners who use GNN training systems and to systems researchers who design and evaluate them.
An extended version of this paper, which includes additional experimental results that we had to omit due to space constraints, is available at (Bajaj et al., 2024). To ease reproducibility, our study uses publicly available datasets, model implementations, and training systems. We make all our scripts publicly available. [^1] https://github.com/saurabhbajaj123/GNN_minibatch_vs_fullbatch
2. Background and Motivation
This section provides necessary background on GNN and GNN training systems to follow the rest of the paper. It then discusses the existing evidence in the literature.
2.1. Graph Neural Networks
Given a graph $G(V,E)$, where each node $v\in V$ is represented with a feature vector $h_{v}^{0}$ (input feature of vertex $v$) and the edges between any two vertices $u$ and $v$ is $e_{uv}\in E$. A GNN at $l^{th}$ layer performs the following computation to get the representation of any vector,
(1) $h_{v}^{l}=\Phi^{l}(h_{v}^{l-1},\Psi^{l}(h_{u}^{l-1},\forall u\in N\cup{v}),W^{l}),\$
where $h_{v}^{l}$ is the $l^{th}$ layer representation for vertex $v$, $N$ represents the incoming neighbor vertices of vertex $v$, $W^{l}$ is the model weight matrix, $\Psi^{l}$ is any aggregation function and $\Phi^{l}$ is an update function.
This paper considers the three standard GNN models used in previous work: GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2017). GraphSAGE learns the representation of a target node based on the features of the nodes in its k-hop neighborhood. The information from the neighboring nodes is propagated to the target node by performing message passing and aggregation recursively, one hop at a time. GAT (Graph Attention Networks) extend message passing by drawing inspiration from attention architectures in transformer models. It uses an attention score to weigh the incoming message of a node. GCN learns hidden layers representations and encode the graph structure and the features of the nodes by implementing a first-order approximation of spectral convolutions on graphs.
Table 1. Review of literature reporting claims of how mini-batch and full-graph training compare. (Notation: TTA - time to accuracy, ACC - accuracy, FG - Full graph, MB - mini-batch, ET - Epoch time)
Proposes System/paper Main Claims Experimental evidence Year Open Source Platform
Full-graph GNNPipe (Chen et al., 2023) FG lower ET ET 2023 Multi-host multi-GPU
Betty (Yang et al., 2023) FG higher ACC ACC 2023 Single-host single-GPU
HongTu (Wang et al., 2023) MB information loss, MB lower ACC ACC, TTA 2023 Single-host multi-GPU, distributed CPU-only
AdGNN (Song et al., 2023a) FG higher ACC, FG better convergence 2023 Multi-host single-GPU
G3 (Wan et al., 2023a) FG higher ACC, lower ET ACC, ET 2023 Multi-host multi-GPU
AdaQP (Wan et al., 2023b) MB information loss 2023 Mutli-host mult-GPU
NeutronStar (Wang et al., 2022) FG lower ET ET 2022 Multi-host single-GPU
PipeGCN (Wan et al., 2022b) FG higher ACC ACC 2022 Multi-host multi-GPU
BNS-GCN (Wan et al., 2022a) FG higher ACC, lower ET ACC, ET 2022 Multi-host multi-GPU
Sancus (Peng et al., 2022) MB information loss 2022 Multi-host multi-GPU
SAR (Mostafa, 2022) MB noisy gradients 2022 Multi-host CPU
Dorylus (Thorpe et al., 2021) FG higher ACC ACC 2021 Multi-host CPU, Multi-host single-GPU
DistGNN (Md et al., 2021) FG lower ET ET 2021 Multi-host CPU
DGCL (Cai et al., 2021) MB ACC loss 2021 Mutli-host mult-GPU
CAGNET (Tripathy et al., 2020) FG higher ACC 2020 Multi-host multi-GPU
ROC (Jia et al., 2020) FG lower TTA TTA 2020 Multi-host multi-GPU
NeuGraph (Ma et al., 2019) FG convergence guarantee 2019 Single-host multi-GPU
Mini-batch BGL (Liu et al., 2023) MB and FB same performance 2023 Multi-host CPU, Multi-host multi-GPU
NeutronOrch (Ai et al., 2023) FG impractical, limited GPU memory 2023 Single-host multi-GPU
DistDGLv2 (Zheng et al., 2022b) MB higher ACC, lower TTA ACC, TTA 2022 Multi-host CPU, Multi-host multi-GPU
SALIENT (Kaler et al., 2022) MB lower ET, higher ACC ACC, ET 2022 Multi-host multi-GPU
GNNLab (Yang et al., 2022b) FG hard to scale 2022 Single-host multi-GPU
ByteGNN (Zheng et al., 2022a) FG not practical for large graphs 2022 Multi-host CPU
CM-GCN (Zhao et al., 2021) MB lower TTA ACC, TTA 2021 Multi-host CPU
Blocking based (Yao and Li, 2021) FG higher memory and computational complexity 2021 -
DistDGL (Zheng et al., 2020) MB lower TTA 2020 Multi-host CPU, multi-host multi-GPU
PaGraph (Lin et al., 2020) MB and FB same performance 2020 Single-host multi-GPU
Cluster-GCN (Chiang et al., 2019) FG - memory: bad; time per epoch: good; convergence: bad 2019 -
LADIES (Zou et al., 2019) MB higher ACC ACC 2019 -
Other papers Rethinking (Li et al., 2024) FG not scalable ACC 2024 -
RDM (Kurt et al., 2023) FG ACC higher or equal, MB faster convergence ACC, TTA 2023 -
EXACT (Liu et al., 2021) FG - higher ACC for small graphs , MB - higher ACC for large graphs ACC 2021 -
OGB (Hu et al., 2020) MB higher ACC ACC 2020 -
2.2. Full-Graph vs. Mini-Batch Training: Different Systems for Different Pipelines
Full-graph and mini-batch training systems implement different data management pipelines for scalable multi-GPU training, which resulted in the emergence of two separate categories of systems. For a more detailed discussion on distributed GNN training systems, we refer the reader to (Shao et al., 2024; Lin et al., 2023).
Full-graph training systems The key idea behind full-graph training systems is to partition a large graph into multiple subgraphs such that each can fit into a GPU, and then train them in parallel. This requires communication of vertex features across GPUs. Different types of communication strategies have been used in the literature, including broadcast-based (Tripathy et al., 2020), point-to-point based (Md et al., 2021; Demirci et al., 2022), pipeline-based (Wan et al., 2023a; Mostafa, 2022; Md et al., 2021; Wang et al., 2021b), and communication via shared memory (Jia et al., 2020; Ma et al., 2019).
Communication incurs a substantial performance overhead and requires frequent synchronization. To overcome this bottleneck, two main types of optimizations have been proposed: asynchronous communication and sampling. Asynchronous communication allows the use of stale local information to avoid expensive synchronization costs. They can be categorized as epoch-fixed (Md et al., 2021; Wan et al., 2022b) and epoch-adaptive (Thorpe et al., 2021) staleness models. In GNN training, two synchronization points exist which are specific to GNN training: first, embedding aggregation and second, gradient aggregation. Some protocols (Peng et al., 2022; Md et al., 2021) only consider the embedding aggregation stage to support asynchronous aggregation, while others (Wan et al., 2022b; Thorpe et al., 2021) are designed to use historical information in both the embedding aggregation and gradient aggregation stages. Sampling keeps training synchronous but reduces communication cost by exchanging messages only among a subset of the boundary vertices across partitions (Wan et al., 2022a). Synchronous training is preferred in many cases because its results are predictable and reproducible.
Mini-batch training systems Mini-batch training systems rely on the generation of mini-batches to scale GNNs to large graphs and employ graph sampling to reduce the neighborhood explosion problem. Mini-batch training systems have a different set of performance bottlenecks than full-graph training ones. Existing research has focused on optimizing two steps, mini-batch sampling and input feature loading, which are major performance bottlenecks and often have a cost comparable to training. Multiple vertex-wise, layer-wise and subgraph-wise sampling techniques have been proposed (Chen et al., 2018a; Hamilton et al., 2017; Zeng et al., 2019; Zou et al., 2019). It is common to accelerate these algorithms using GPUs (Jangda et al., 2021; Pandey et al., 2020; Wang et al., 2021a; Gong et al., 2023; Yang et al., 2022b). To avoid massive CPU-GPU communication while loading, some works (Zhu et al., 2019) have proposed caching techniques, like replicating frequently accessed remote vertices at local , store high out-degree vertices (Lin et al., 2020), caching hot vertices (i.e., those accessed more often) (Min et al., 2022). Some works (Liu et al., 2023) propose dynamic caching using a FIFO policy. Some distributed sampling techniques (Jiang and Rumi, 2021; Cai et al., 2023) prioritize local vertices, pushing sampling tasks to remote workers only if results need to be sent back, reducing communication. Some cache both topological and feature data on the GPUs (Sun et al., 2023; Zhang et al., 2023; Cai et al., 2023). To avoid graph sampling becoming a bottleneck, some works have focused on developing efficient GNN dataloaders (Bai et al., 2021; Deutsch et al., 2019; Dong et al., 2021; Min et al., 2021). Some systems reduce the cost of loading by proposing hybrid parallelism approaches for mini-batch training that are alternative to data parallelism: push-pull parallelism (Gandhi and Iyer, 2021) and split parallelism (Polisetty et al., 2023).
2.3. Motivation
The motivation for this work is that prior work on GNN training systems makes it difficult to establish the state of the art in the field. A significant amount of works on GNN training do not compare or evaluate their systems against systems using a different training approach(full-graph or mini-batch) (Kaler et al., 2023; Sun et al., 2023; Cai et al., 2023; Chen et al., 2018a; Hamilton et al., 2017; Chen et al., 2018b; Huang et al., 2018; Zeng et al., 2019; Gasteiger et al., 2022; Wang et al., 2021a; Zhu et al., 2019; Zhang et al., 2020; Kipf and Welling, 2017; Gandhi and Iyer, 2021). In the following, we review 33 papers on GNN training that do make claims comparing systems in the two classes. We summarize the claims in these papers in Table 1.
We classify the papers into three broad groups. The first group consists of papers that propose full-graph training. Out of the 17 papers in this group, 10 provide experimental evidence. The rest 7 papers make claims that full-graph training is superior in terms of higher accuracy, lower epoch time, and convergence guarantees and support the claims by only citing other papers. The second group includes 12 papers that propose mini-batch training. Among these papers, only 4 provide experimental evidence to support claims that full-graph training is difficult to scale and requires more memory, while mini-batch training has lower time-to-accuracy. The last group consists of 4 papers that do not propose one specific training approach. Overall, we find the following shortcomings in the existing literature, which motivate our comprehensive experimental comparison.
Conflicting claims about how the two approaches compare. Several papers proposing full-graph systems (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wang et al., 2023; Song et al., 2023a) demonstrate that their systems achieve higher accuracy and lower epoch time compared to mini-batch systems. Some papers mention that mini-batch causes information loss, noisy gradients, and does not guarantee convergence (Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019; Wang et al., 2023; Song et al., 2023a) However, papers proposing mini-batch systems (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020) and some other works (Hu et al., 2020) present contradicting evidence and claims showing that mini-batch training leads to higher accuracy and faster convergence to the target accuracy level. Other works claim full-graph training is impractical, requires high memory and is hard to scale (Ai et al., 2023; Yang et al., 2022b; Zheng et al., 2022a; Chiang et al., 2019; Yao and Li, 2021). Some argue that no approach is clearly superior and assessing their effectiveness requires a comprehensive evaluation which is out of scope of their work (Liu et al., 2021; Kurt et al., 2023; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Liu et al., 2023; Lin et al., 2020; Wang et al., 2023).
Comparing performance only in terms of epoch time. Certain works compare full-graph and mini-batch training systems based only on epoch time (Wang et al., 2022; Wan et al., 2023a; Md et al., 2021; Chen et al., 2023). As we will show, a comparison in terms of time to accuracy is more informative given the differences in the two training approaches. Our observations indicate that although mini-batch training exhibits higher epoch time, it still takes less time to reach a target accuracy level (Section 4.1).
Experiment design. Some works (Jia et al., 2020; Liu et al., 2021; Li et al., 2024; Kaler et al., 2022) do not always use the same GNN model (e.g., GraphSAGE or GAT) for the two approaches as a basis of comparison. This conflates the effect of the training approach with the GNN model choice, since different models may exhibit varying performance even when trained with the same technique. Some papers do not mention explicitly whether they perform separate hyperparameter tuning for each method (Wan et al., 2023a, 2022b, 2022a; Thorpe et al., 2021; Zhao et al., 2021; Kaler et al., 2022). As we show in our experimental evaluation of Section 5, given a dataset and GNN model, it is possible to find hyperparameter settings where one method achieves higher accuracy than another and vice versa. Using hyperparameter tuning is necessary to perform a fair comparison.
Table 2. Dataset statistics. (*) The input features and the training data for Orkut are synthetic
Datasets #Nodes #Edges #Classes #Features Train/Val/Test
Pubmed 19,717 108,365 3 500 18,217/500/1,000
Ogbn-arxiv 169,343 1,166,243 40 128 90,941/29,799/48,603
Reddit 232, 965 11, 606, 919 41 602 152,410/23,699/55,334
Ogbn-products 2,449,029 61,859,140 47 100 195,922/48,980/2,204,127
Orkut 3,072,441 117,185,083 2* 128* 1,843,464/614,488/614489*
Ogbn-papers100M 111,059,956 1,615,685,872 172 128 1,207,179/125,265/214,338
Table 3. Hyperparameter search space
Parameter range/values
Number of Hidden Layers [2, 10]
Size of hidden layers [16, 1024]
Aggregator (GraphSage only) { mean, gcn, pool }
Number of heads (GAT only) [1, 12]
Sampling Fanout (MB only) [4, 50]
Learning rate [0.0001, 0.01]
Dropout [0.2, 0.9]
Mini-batch size (MB only) [128, 4096]
3. Experimental setup
Hardware setup. We run our experiments using two types of hosts. The first is a lower-end one, which we call PCIe. It has 2 Intel Xeon E5-2620 v3 CPUs with 12 cores each, 256GB of host memory, and 4 NVIDIA Tesla m40 GPUs, each having 24GB of memory, connected via PCIe. These servers are connected with a 1 GBps network. The second host type is higher end and we call it NVLink. It has 2 Intel Xeon Platinum 8480+ CPUs with 56 cores each, 256 GB of host memory, and 4 A100 GPUs, each with 80GB of memory. These servers are connected with a 25 GBps network. By default, all our experiments use 4 GPUs per host.
Datasets. Our experiments use six datasets: pubmed (Namata et al., 2012), ogbn-arxiv (Hu et al., 2020), reddit (Hamilton et al., 2017), ogbn-products (Hu et al., 2020), orkut (Yang and Leskovec, 2012) and ogbn-papers100m (Hu et al., 2020) (see Table 2). Except orkut, all these datasets come with input features that enable evaluating the accuracy of GNN training. The orkut dataset contains only the graph topology, so we only use it to measure epoch time for the scalability experiments. To run GNN training on it, we add synthetic input features of size 128, a train/val/test split of 60%/20%/20%. One out of two target classes is randomly assigned to each node.
Representative GNN training systems. To compare across full-graph and mini-batch distributed GNN systems, we select representative systems that incorporate the main optimizations discussed in Section 2.2. We chose systems that have a stable open-source implementation, including implementations of our target models (GCN, GraphSage, and GAT) in their repositories [^2] DGL 1.1.2: https://github.com/dmlc/dgl; Quiver 0.1.0: https://github.com/quiver-team/torch-quiver; DistDGL: https://github.com/dmlc/dgl/tree/master/examples/distributed; PipeGCN: https://github.com/GATECH-EIC/PipeGCN; BNS-GCN: https://github.com/GATECH-EIC/BNS-GCN. We run all systems on top of PyTorch 2.0.1 and Python 3.8.10.
For full-graph training, we select PipeGCN (Wan et al., 2022b), which supports asynchronous training to overlap communication and computation, and BNS-GCN (Wan et al., 2022a), which uses sampling to reduce communication. Both systems partition the entire graph topology and features, and store each partition in the memory of one GPU.
For mini-batch training, we choose popular data-parallel systems that support GPU-based sampling, efficient input feature loading, and feature caching. Specifically, we use DGL (Wang et al., 2019), which only supports single-GPU caching, and Quiver (Tan et al., 2023), which supports distributed caching. DistDGL (Zheng et al., 2020) is built on top of DGL and performs mini-batch training on a distributed cluster of hosts. It only supports CPU-based training and does not support caching. All the mini-batch systems we consider support only synchronous training.
Model selection and implementations. We consider standard GNN models that are commonly used to evaluate training systems: GCN, GraphSage, and GAT. We use the implementations provided with the distributions of the systems we benchmark and validate that the semantics of the implementations are consistent across different systems. This is to make sure that the model implementations are fully compliant and optimized for the respective systems and avoid introducing biases. DGL and Quiver include implementations of all the three models in its distribution. The PipeGCN distribution includes GraphSAGE, while the GCN and GAT implementations for both PipeGCN and BNS-GCN are included in the BNS-GCN distribution. For mini-batch training, we considered the standard Neighborhood Sampling algorithm for our evaluation (Hamilton et al., 2017).
Hyperparameter search space. The search space for hyperparameters is shown in Table 3. The first set of hyperparameters define the architecture of the GNN models. Fanout is related to the neighborhood sampling algorithm used in mini-batch training. Finally, we consider general model training parameters such as learning rate, dropout and batch size. To search this parameter space, we use random search for all the parameters to obtain the best set of hyperparameters giving the highest validation accuracy.
4. Performance Evaluation
We evaluate the performance of our representative GNN training systems and answer two questions: Q1) How long do representative systems take to reach the same target accuracy? Q2) How do these systems scale with a varying number of GPUs and hosts? We mainly use time-to-accuracy as the main metric for performance comparison since it is a more informative metric than epoch time, as we will discuss.
We conduct end-to-end performance measurements because mini-batch and full-graph training systems have different implementations and use different optimizations. For a more in-depth discussion of the effect of specific systems optimizations, it is necessary to restrict the evaluation to one class of systems, as done in (Yuan et al., 2024) for mini-batch systems.
4.1. Time-to-Accuracy
We start by answering question Q1: How long do representative systems take to reach the same target accuracy?
Methodology Time-to-accuracy is the time taken by any system to train to a target accuracy. To measure it, we select the target accuracies by leveraging hyperparameter tuning. Given a dataset and a GNN model, we first find a model architecture on which both training approaches (full-graph and mini-batch) converge to a similar and close-to-best accuracy, which we call convergence accuracy. We select the target accuracy as the minimum convergence accuracy across the two training approaches. We then use the same model architecture across all GNN training systems and measure their time to reach the target accuracy.
These experiments consider the two host types described in Section 3: a higher-end NVLink host and a lower-end PCIe host.
4.1.1. Single NVLink host
Table 4. Time to accuracy for GraphSAGE on a single NVLink host with 4 GPUs. Architectures (# layers, hidden size): pubmed - (3,256), ogbn-arxiv - (2,512), reddit - (3,1024), ogbn-products - (3,256), ogbn-papers100m - (2, 128). ET - Epoch time, TTA - Time to accuracy. For ogbn-papers100m DGL uses UVA mode for sampling and Quiver performs distributed caching. (Best results shown in bold.)
System pubmed ogbn-arxiv reddit ogbn-products ogbn-papers100m
Epoch time (s) PipeGCN (FG, Sync) 0.009 0.057 1.845 2.086 1.094
PipeGCN (FG, Async) 0.008 0.011 0.370 0.341 0.644
BNS-GCN (FG, Sync) 0.012 0.017 0.483 0.473 0.846
DGL (MB, Sync) 0.007 0.110 2.262 1.335 3.428
Quiver (MB, Sync) 0.005 0.131 1.994 2.117 1.878
Time to accuracy (s) PipeGCN (FG, Sync) 0.87 11.03 202.95 479.81 984.33
PipeGCN (FG, Async) 1.02 3.44 103.47 122.64 553.75
BNS-GCN (FG, Sync) 1.59 4.51 144.98 217.66 262.11
DGL (MB, Sync) 0.44 1.10 6.78 46.73 75.42
Quiver (MB, Sync) 0.36 0.65 9.97 84.70 41.31
Table 5. Time to accuracy with 3 NVLink hosts using the ogbn-papers100m graph and a GraphSAGE model with 2 layers and a hidden size of 128 (DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached). (Best results shown in bold.)
System Time (s)
Epoch time (s) PipeGCN (FG, Sync) 1.139
PipeGCN (FG, Async) 0.352
BNS-GCN (FG, Sync) 0.432
DistDGL (MB, Sync) 4.452
Time to accuracy (s) PipeGCN (FG, Sync) 1093.63
PipeGCN (FG, Async) 306.59
BNS-GCN (FG, Sync) 133.92
DistDGL (MB, Sync) 35.62
Table 6. Time to accuracy and epoch time for GraphSAGE with a single PCIe host. Notation: ET - Epoch time, TTA - Time to accuracy. Architectures (# layers, hidden size): pubmed (3,256), ogbn-arxiv (2,512), reddit (4,1024) & ogbn-products (5,256). (Best results shown in bold.)
System pubmed ogbn-arxiv reddit ogbn-products
ET (s) PipeGCN (FG, Sync) 0.024 0.188 5.602 5.835
PipeGCN (FG, Async) 0.020 0.067 3.278 2.526
BNS-GCN (FG, Sync) 0.031 0.072 3.399 2.915
DGL (MB, Sync) 0.019 0.314 9.416 7.081
Quiver (MB, Sync) 0.019 0.216 10.272 9.246
TTA (s) PipeGCN (FG, Sync) 2.40 36.16 616.25 1341.95
PipeGCN (FG, Async) 2.35 20.75 917.80 909.43
BNS-GCN (FG, Sync) 2.91 19.33 1019.64 1340.80
DGL (MB, Sync) 1.25 3.14 28.25 247.83
Quiver (MB, Sync) 1.32 1.08 51.36 369.85
Table 7. Time to accuracy and epoch time with 3 PCIe hosts using the ogbn-papers100m graph and a GraphSAGE model with 2 layers and a hidden size of 128 (DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached). (Best results shown in bold.)
System Time (s)
Epoch time (s) PipeGCN (FG, Sync) 22.00
PipeGCN (FG, Async) 1.03
BNS-GCN (FG, Sync) 2.46
DistDGL (MB, Sync) 38.37
Time to accuracy (s) PipeGCN (FG, Sync) 19141.48
PipeGCN (FG, Async) 956.34
BNS-GCN (FG, Sync) 859.26
DistDGL (MB, Sync) 767.38
Figure 2. Time to accuracy for ogbn-arxiv, reddit, ogbn-products for GraphSAGE (single NVLink host).
Figure 3. Time to accuracy for ogbn-papers100M on (3 NVLink hosts)
Table 4 shows the time to accuracy when training GraphSAGE on a single NVLink host. Full-graph training systems have a consistently lower epoch time (ET) than mini-batch training systems because they perform only one training iteration per epoch. However, their time-to-accuracy (TTA) is substantially higher compared to mini-batch systems. For example, on the largest graph we consider, ogbn-papers100m, Quiver is 6.4$\times$ faster than BNS-GCN. Similarly, for pubmed, Quiver is 2.4×2.4$\times$2.4 × faster than PipeGCN sync (lowest speed-up we observed) and for reddit, DGL is 15.2×15.2$\times$15.2 × faster than PipeGCN async (highest speed-up we observed). On average, the fastest mini-batch systems are 6.3×6.3$\times$6.3 × faster than the fastest full-graph system we considered for these datasets. This is because mini-batch training systems perform multiple training iterations and model updates per epoch and thus converge in fewer epochs.
Figure 2 shows the time-to-accuracy curves. The dotted line indicates the target accuracy. We observe mini-batch training consistently has a much faster time to accuracy across graphs with different sizes. Results on other models are shown in the extended version (Bajaj et al., 2024).
The experiments also show the impact of the systems optimizations introduced in previous work. Among the full-graph training systems, PipeGCN can substantially decrease epoch time and time-to-accuracy using asynchronous training compared to its baseline synchronous version. Sampling, as proposed by BNS-GCN, is less effective than asynchronous training on the smaller graphs, but it yields the fastest time-to-accuracy among all full-graph training systems on ogbn-papers100m.
All full-graph systems we consider fully cache the graph in GPU memory. Mini-batch training systems also cache all the graph in the GPU memory with the exception of ogbn-papers100m, which cannot be cached by DGL. This is because DGL only supports caching when the entire graph fits in the memory of each GPU, unlike Quiver which implements a distributed cache where the graph can be partitioned among multiple GPUs. Despite the lack of caching, DGL still achieves a lower time-to-accuracy than the full-graph training systems on ogbn-paper100m.
4.1.2. Multiple NVLink hosts
We perform a similar experiment using three NVLink hosts with the largest of the graphs we consider, ogbn-papers100m.
Table 5 shows the results. Figure 3 shows the corresponding time-to-accuracy curves. Quiver does not support multi-node training, so we do not include it in these experiments. We also replace DGL with its distributed version, DistDGL. The main trends are similar to the ones observed for the single-host experiments. DistDGL, which uses mini-batch training, has a much larger epoch time than the other systems, which use full-batch training. Its time-to-accuracy, however, is much shorter than the other systems. It is worth noting that DistDGL has a much larger time-to-accuracy than DGL on a single host (see Table 4). This is because DistDGL uses CPU-based sampling and does not support caching, unlike DGL.
4.1.3. PCIe hosts
Table 6 reports the results for single-host training with GraphSAGE using a lower-end host setup, where NVLink is not available and cross-GPU communication within one host must occur over a slower PCIe bus. Compared to the NVLink experiments of Table 4, the epoch time of all systems is larger because the GPUs are less powerful. Full-graph training systems experience a stronger degradation in performance than DGL because they must run cross-GPU communication over PCIe, whereas DGL avoids communication by using data parallelism. Quiver also observes a performance degradation since loading data from other GPU’s cache is not as effective over the PCIe bus as it is over NVLink.
Table 7 futher compares all systems using 3 PCIe hosts for the ogbn-papers100m graph. In the PCIe setup, full-graph training systems need a multi-node setup to fully cache the ogbn-papers100m graph, which is larger than the others. In the setup with 3 PCIe hosts, hosts are connected by a slower network (1 GBps instead of 25 GBps). Like in the previous experiments, DistDGL has a lower time-to-accuracy than the other full-graph training systems, but the difference is much smaller than with the NVLink setup. This is because DistDGL requires cross-host communication to perform sampling and feature loading. These costs become bottlenecks over the slow 1GBps network of the PCIe host. By contrast, full-graph training systems are optimized to minimize communication and amortize its cost.
Takeaway. Mini-batch training enables faster time-to-accuracy compared to full-graph training since it requires fewer epochs to reach a target accuracy. This is because full-graph training systems update the model only once per epoch, whereas mini-batch training systems perform multiple iterations per epoch and update the model at each iteration. By updating the model more frequently, mini-batch training systems typically converge faster than full-graph training ones, even if their epoch time is longer.
These findings stress that time-to-accuracy measurements should always be included to provide a comprehensive comparison of systems using different training approaches. Comparing only in terms of epoch time, as is commonly done in prior work (see Table 1) does not provide a comprehensive comparison.
4.2. Scalability
Table 8. Single-host scalability with the GAT model using the NVLink host. Times in seconds. (#Layers, hidden size, #heads, fanout): ogbn-arxiv - (3, 1024, 6, 25), Reddit - (2, 1024, 4, 20), ogbn-products - (3, 128, 3, 20), Orkut - (3, 128, 2, 5). ET - Epoch time, TTA - Time to Accuracy.
Dataset ogbn-arxiv reddit ogbn-products orkut
System # of GPUs ET TTA TTA Speed-up ET TTA TTA Speed-up ET TTA TTA Speed-up ET ET speed-up
PipeGCN (Sync) 1 1.293 569.01 1.0x 4.265 4094.59 1.0x 2.058 823.36 1.0x 1.898 1.0x
2 0.868 381.74 1.5x 2.566 2463.65 1.7x 1.341 536.20 1.5x 1.873 1.0x
3 0.763 335.90 1.7x 2.587 2483.90 1.6x 1.129 451.64 1.8x 1.594 1.2x
4 0.625 275.04 2.1x 1.801 1728.58 2.4x 1.033 413.28 2.0x 1.442 1.3x
PipeGCN (Async) 1 1.292 581.27 1.0x 4.268 4267.70 1.0x 2.060 782.76 1.0x 1.898 1.0x
2 0.777 349.83 1.7x 2.309 2309.20 1.8x 1.135 431.34 1.8x 1.149 1.7x
3 0.589 265.10 2.2x 2.309 2309.20 1.8x 0.810 307.61 2.5x 0.920 2.1x
4 0.469 211.14 2.8x 1.312 1312.15 3.3x 0.614 233.24 3.4x 0.720 2.6x
BNS-GCN (Sync) 1 1.309 471.06 1.0x 4.189 4188.60 1.0x 2.171 455.95 1.0x 2.173 1.0x
2 0.764 274.93 1.7x 2.315 2315.30 1.8x 1.213 254.73 1.8x 1.284 1.7x
3 0.532 191.56 2.5x 1.724 1723.60 2.4x 0.857 179.99 2.5x 0.984 2.2x
4 0.420 151.20 3.1x 1.395 1395.40 3.0x 0.629 132.13 3.5x 0.713 3.0x
DGL (Sync) 1 15.006 150.06 1.0x 24.006 360.09 1.0x 10.554 158.31 1.0x 14.791 1.0x
2 7.290 72.90 2.1x 12.794 191.91 1.9x 5.827 87.40 1.8x 7.668 1.9x
3 5.550 61.04 2.5x 8.949 134.24 2.7x 3.894 58.42 2.7x 4.940 3.0x
4 3.871 42.58 3.5x 6.510 97.65 3.7x 2.730 40.95 3.9x 3.980 3.7x
Quiver (Sync) 1 13.813 138.13 1.0x 26.427 396.41 1.0x 11.212 168.19 1.0x 13.718 1.0x
2 6.575 65.75 2.1x 13.764 206.46 1.9x 5.797 86.95 1.9x 7.447 1.8x
3 4.979 49.79 2.8x 8.703 130.55 3.0x 3.993 59.90 2.8x 5.058 2.7x
4 3.596 35.96 3.8x 6.779 101.68 3.9x 2.974 44.61 3.8x 3.855 3.6x
Table 9. Distributed scalability with the GraphSage model using 4 GPUs per host connected with NVLink . Times in seconds. (#Layers, hidden size, fanout): Orkut - (3, 512, 10), ogbn-products100M - (2, 512, 10). ET - Epoch time, TTA - Time to Accuracy. DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached
orkut ogbn-papers100m
System # Hosts ET ET Speed-up ET TTA TTA Speed-up
PipeGCN 1 4.769 1.0x 3.556 924.46 1.0x
(Sync) 2 4.588 1.0x 3.566 927.13 1.0x
3 4.351 1.0x 3.933 1022.63 0.9x
PipeGCN 1 0.454 1.0x 0.759 197.57 1.0x
(Async) 2 0.341 1.3x 0.464 120.67 1.6x
3 0.227 2.0x 0.361 93.96 2.1x
BNS-GCN 1 0.810 1.0x 0.997 209.38 1.0x
(Sync) 2 0.608 1.3x 0.762 160.10 1.3x
3 0.405 2.0x 0.542 113.88 1.8x
DistDGL 1 9.826 1.0x 11.561 80.93 1.0x
(Sync) 2 7.305 1.3x 10.749 75.24 1.1x
3 6.815 2.0x 6.612 46.28 1.7x
Next we answer Q2: How do systems scale with a varying number of GPUs and a varying number of hosts? For each dataset, we pick a GNN architecture that fits in a single GPU and use the single-GPU training performance as a basis to calculate the scalability speedups.
4.2.1. Single-host scalability (NVLink)
We now report the single-host scalability of our two approaches by varying the number of GPUs. We consider the GAT model since it is the most computationally demanding among the ones we considered.
Table 8 presents the results from our scalability experiment. The performance trends when training on a single GPU are similar to the ones we observed for the multi-GPU setups we discussed previously. The full-graph training systems have a much lower epoch time than the mini-batch ones but a higher time-to-accuracy.
When scaling to multiple GPUs, one important advantage of mini-batch training systems is that they can keep the amount of work per GPU constant by increasing the mini-batch size. This is not possible in full-graph training systems since the total amount of work per epoch is constant. As the number of GPU increases, the amount of work per GPU decreases but the communication cost increases, making it difficult to achieve scalability.
This effect is clear when comparing systems using synchronous training. With PipeGCN, which uses full-graph training, the cost of synchronous cross-GPU communication during the forward and backward pass of training hides the gains of reducing the computation cost per GPU.
Several optimizations in full-graph training systems address this problem. PipeGCN has a much better scalability when using asynchronous training. This enables PipeGCN to use pipelining, one of its most important features, to overlap communication and computation and to scale to up to 3.4×3.4$\times$3.4 ×. Sampling of boundary nodes in BNS-GCN also enables good scaling, up to 3.5×3.5$\times$3.5 ×, since the cross-GPU communication is avoided. The mini-batch systems scale almost linearly despite using synchronous training: both DGL and Quiver scale by up to 3.9×3.9$\times$3.9 ×. Besides scalability, these systems achieve lower time-to-accuracy that the full-graph training systems in all settings. We only show ET and ET Speed-up for orkut dataset, since it is a SNAP dataset (Yang and Leskovec, 2012) with only graph structure and synthetic input features and target labels.
We also experiment with GraphSAGE model, and the results are shown in the extended version (Bajaj et al., 2024). For GraphSAGE we see that synchronous PipeGCN doesn’t scale at all, asynchronous PipeGCN scales up to $\times3$ and BNS-GCN scales up to 2.7×2.7$\times$2.7 ×. While DGL and Quiver are able to scale up to 4.0$\approx$ 4.0$\times$ 4.0 × and 3.8×3.8$\times$3.8 × respectively.
Takeaway. The mini-batch training systems show better scalability than the full-graph training ones. This is because they can tune the mini-batch size to keep the work per GPU constant to maintain high GPU utilization, unlike the full-graph training systems. The full-graph training systems also experience increasing communication costs as the number of GPUs increases. However, these overheads can be mitigated using asynchronous training and boundary nodes sampling.
4.2.2. Distributed scalability (NVLink)
We now consider scaling the training of larger graphs (orkut and ogbn-papers100m) in a distributed setting of NVLink hosts, where each host has four GPUs. We consider the GraphSAGE model because it is the only one implemented in the DistDGL distribution. We do not consider Quiver because it does not support distributed training.
Table 9 shows the results for distributed scalability with NVLink enabled devices. The synchronous version of PipeGCN cannot scale when increasing the number of hosts. Using asynchronous training and boundary nodes sampling can make full-graph training systems scale much better, as we also observed for the single-host experiments. DistDGL has similar scalability trends despite the fact that it does not use GPU caching. Despite its higher epoch time compared to full-graph training systems, DistDGL achieves a much lower time-to-accuracy.
Takeaway. The general scalability considerations we made for single-host scalability apply also to multi-host scalability. Synchronous full-graph training does not scale to multiple host. By using asynchronous training and boundary nodes sampling, PipeGCN and BNS-GCN achieve similar scalability as DistDGL, which uses mini-batch training. However, DistDGL has a lower time-to-accuracy in all configurations.
4.2.3. Single-host scalability (PCIe hosts)
We run the same experiments on PCIe hosts (as done in Section 4.1.3), and observe similar results, which are shown in the. We scale the number of GPUs from one to four.
All the full-graph training systems have a lower scalability because they require communication over the PCIe bus, which is slower than NVLink. This slowdown compared to the NVLink case is particularly evident for PipeGCN sync, which can only scale up to 2.0×2.0$\times$2.0 ×. PipeGCN async can scale up to at most 3.1×3.1$\times$3.1 × and BNS-GCN up to at most 3.0×3.0$\times$3.0 ×.
Mini-batch training systems have a better scalability because they require little cross-GPU communication. DGL scales up to 3.8×3.8$\times$3.8 × and Quiver up to 3.5×3.5$\times$3.5 ×.
Takeaway. The full-graph systems are more sensitive to the performance of the cross-GPU interconnect because they need to exchange vertex features between partitions at each GNN layer during the forward and backward pass. When GPUs must communicate over a slow PCIe interconnect, the scalability of these systems is lower. The mini-batch training systems use data-parallelism, which does not require exchanging vertex features across GPUs. Therefore, these systems show better scalability in the PCIe host.
5. Accuracy Evaluation
Even though a system takes longer than another to reach the same target accuracy, it could still be preferable to use it if it can converge to a higher accuracy. Indeed, some works on GNN systems assert that full-graph training achieves higher accuracy and that mini-batch training does not guarantee convergence (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
Figure 4 summarizes the accuracy values reported in the literature we reviewed for different GNN models and datasets. Each point in the graph represents the test accuracy reported for a specific GNN model-dataset pair in one of the papers we reviewed. The figure illustrates a wide distribution of accuracy values for the same dataset and model combination.
In this Section, we shed light onto this apparent contradiction and answer two questions: Q3) What is the best accuracy we can achieve using full-graph and mini-batch training? Q4) What accuracy can we achieve when using a training method with the best hyperparameters found for the other method? In practice, the cost of training a model also depends on how long it takes to find good hyperparameters. Therefore, we also consider the following question: Q5) How sensitive are the two training methods to changes in the hyperparameters?
Figure 4. Test accuracy reported in literature on GNN training. (Wan et al., 2022b; dgl, 2023; Hu et al., 2020; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Md et al., 2021; Yang et al., 2023; Li et al., 2024; Gasteiger et al., 2022; Mostafa, 2022; Kaler et al., 2022; Yao and Li, 2021; Zou et al., 2019; Hamilton et al., 2017; Chiang et al., 2019; Zeng et al., 2019; Chen et al., 2018b, a; Yao and Li, 2021; Wu et al., 2024; Veličković et al., 2018; Thorpe et al., 2021; Frasca et al., 2020; Wan et al., 2022a; Kipf and Welling, 2017; Jia et al., 2020)
Methodology To compare maximum accuracies among full-graph and mini-batch training, we use the methodology depicted in Figure 5. Given a dataset and a GNN model, we conduct two separate hyperparameter searches, one using full-graph training and one using mini-batch training. This is because the two training methods can have different sets of best hyperparameters for the same dataset and model. The space of hyperparameters we consider is shown in Table 3. The best hyperparameter settings we obtained using the two training methods are denoted as FG-search and MB-search. We then run full-graph and mini-batch training on both settings, measure the test accuracy, and denote it as FG-train and MB-train respectively. This yields a total of four accuracy combinations per dataset and GNN model. We report all the hyperparameter settings we found in the extended version of this paper (Bajaj et al., 2024).
To factor out any accuracy bias due to systems-level optimizations, we perform synchronous training with PipeGCN and DGL. We validate our results by verifying that given a dataset, model, and training method, our hyperparameter search achieves accuracy values that are equal to or better than the best accuracy value reported in the literature reviewed in Section 2.3.
Figure 5. Methodology of our accuracy comparison between full-graph (FG) and mini-batch (MB) training.
5.1. Accuracy with the Best Setting
We start by answering question Q3) What is the best accuracy we can achieve using full-graph and mini-batch training? Answering this question shows whether there is one method that consistently reach higher accuracy than another. When the hyperparameters are properly set, whether full-graph training achieves higher accuracy than mini-batch training or vice versa depends on the dataset-GNN model pairs.
The accuracy results for the GraphSAGE model are shown in Table 10. Given a dataset, we must compare the accuracies obtained for two combinations: (FG-search, FG-train) for full-graph training and (MB-search, MB-train) for mini-batch training. The results are averaged after 5 runs.
The results show that the accuracies are similar across training methods. For the GraphSAGE model, each method can yield higher accuracy than the other, depending on the dataset. The largest gap between the accuracies obtained with the two methods is less than $0.9\%$ in favor of full-graph training and averaging around $0.3\%$. We ran a similar analysis for the GCN and GAT model, and we report the results in the extended version of this paper (Bajaj et al., 2024). With the GAT model, the largest gap is slightly higher, up to $1.1\%$ in favor of full-graph training and the average is around $0.6\%$. Gaps are higher for GCN, which is a simpler model and turns out to be more sensitive to the choice of hyperparameters. The largest gap in this case is $3.6\%$ in favor of mini-batch training and the average is around $1.27\%$.
Takeaway. These results indicate that no training method alone guarantees higher accuracy than the other and the gap between the two methods is typically small (i.e., less than 3.6% accuracy). If achieving the highest accuracy is an absolute priority regardless of the training cost, it is advisable to conduct hyperparameter search twice using both training methods. Otherwise, using either of the two methods is likely to produce similar accuracy in the absence of other systems-level optimizations such as asynchronous training or boundary node sampling.
Table 10. Accuracy (%) of GraphSage using hyperparameter tuning method in Figure 5.
Pubmed Ogbn-arxiv Reddit Ogbn-products Ogbn-papers100M
FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search
FG-train $\textbf{77.40}\pm\textbf{00.04}$ $76.10\pm 00.67$ $\textbf{71.76}\pm\textbf{00.16}$ $67.21\pm 00.9$ $\textbf{96.94}\pm\textbf{00.01}$ $90.28\pm 00.63$ $\textbf{79.29}\pm\textbf{00.15}$ $76.60\pm 00.32$ $\textbf{62.82}\pm\textbf{00.01}$ $61.74\pm 00.13$
MB-train $77.00\pm 00.39$ $\textbf{78.20}\pm\textbf{00.32}$ $70.73\pm 00.38$ $\textbf{71.67}\pm\textbf{00.30}$ $96.52\pm 00.14$ $\textbf{96.82}\pm\textbf{00.04}$ $78.72\pm 00.07$ $\textbf{78.84}\pm\textbf{00.12}$ $62.52\pm 00.04$ $\textbf{62.77}\pm\textbf{00.02}$
5.2. Accuracy with Sub-Optimal Setting
We now answer the next question Q4) What accuracy can we achieve when using a training method with the best hyperparameters found for the other method? This shows how the outcome of the comparison between the two training approaches can change when using sub-optimal hyperparameter settings.
To answer this question for the case of the GraphSAGE model, we still look at Table 10. Given a dataset, the FG-search column reports the accuracy achieved with the two training methods using the hyperparameter setting where full-graph training works best. Mini-batch training performs worse than full-graph training, but only by a small margin. The largest gap between MB-train and FG-train is for the ogbn-arxiv graph ($1.03\%$).
The MB-search column compares the two training methods using the best hyperparameters we found for mini-batch training. Full-graph training has lower accuracy than mini-batch training and the gap between the two methods is larger, reaching $6.22\%$ for the reddit dataset.
Our evaluation on GAT and GCN, which is reported on the extended version (Bajaj et al., 2024), leads to similar conclusions. For GAT, largest difference between FG-train and MB-train is $2.85\%$ in favor of MB-search for the ogbn-arxiv dataset and $1.9\%$ in favor of FG-search for the pubmed dataset. For GCN, it is $1.36\%$ in favor of FG-search for the reddit dataset and $5.5\%$ in favor of MB-search for the pubmed dataset.
Takeaway. These results show that given a dataset and GNN model, each training method can outperform the other depending on the choice of the hyperparameters. This corroborates our choice of performing two separate hyperparameter searches, one per training method, for our comparison. It also potentially explains why existing work shows inconsistent evidence about which training method leads to higher accuracy. We hypothesize that the large variations in accuracy shown in Figure 4 is due to the difference in the choice of hyperparameter settings, which can greatly impact the accuracy and performance of GNN training.
Mini-batch training still performs relatively well in settings where full-graph training performs best, but the opposite is not true. This is an indication that mini-batch training is more robust to different hyperparameter settings, potentially making hyperparameter search easier. We will further corroborate this thesis in later experiments.
5.3. Ease of Hyperparameter Search
Figure 6. Sensitivity to the number of layers and hidden size for GraphSAGE model.
We finally answer our last question focusing on accuracy and hyperparameter search, which is Q5) How sensitive are the two training methods to changes in the hyperparameters? We evaluate the ease of finding a good hyperparameter setting using the two training approaches by measuring how accuracy varies as we vary the model architecture hyperparameters, i.e., the number of hidden layers and the size of the hidden vertex dimensions.
We trained each model for a maximum of 1000 epochs, stopping earlier if we the validation accuracy doesn’t increase for 50 epochs.
Figure 6 illustrates the best test accuracy when training the GraphSAGE model, top row shows mini-batch training experiments and the bottom row shows synchronous full-graph training. Notably, mini-batch training demonstrates remarkable consistency across varying model architectures. In contrast, full-graph training exhibits a wider range of variations in test accuracy when the architecture changes. We observed similar results for GCN and GAT models, which are shown in the extended version of this paper (Bajaj et al., 2024).
Figure 6 and the other results for GCN and GAT, which are reported in the extended version (Bajaj et al., 2024), clearly show the effect of over-smoothing, which appears in GNNs at increased depths (Rusch et al., 2023). Mini-batch training on GraphSAGE does not show smoothing with up to 8 layers in Figure 6. In a set of separate experiments we found that oversmoothing does occur with mini-batch training, but only after we increase the number of layers to $16$ or higher. Our experiments for GCN and GAT show that oversmoothing occurs at lower depths for mini-batch training but the resulting accuracy drop is still much less than that with full-graph training (see the extended version (Bajaj et al., 2024)).
Takeaway. These results highlight the robustness of mini-batch training to alterations in model architecture. Full-graph training appears to be more sensitive to changes in architecture, so it requires a more careful and potentially expensive hyperparameter search. Whereas mini-batch training is consistently less sensitive to changes of those parameters than full-graph training, across all the GNN models we consider. In addition, mini-batch training is much less prone to the oversmoothing problem (Rusch et al., 2023) that affects deep GNN models with many layers.
6. Conclusion
This work is motivated by the need to understand the state of the art in GNN training systems, given the absence of consistent empirical evidence comparing full-graph and mini-batch training systems. It first conducted an in-depth literature survey and investigated the reasons for the inconsistent claims and the lack of empirical evidence in the literature. It further provided a thorough empirical evaluation comparing representative systems supporting the two approaches on various GNNs and datasets.
The study reveals that the mini-batch training systems we considered consistently achieve faster convergence time, measured as time-to-accuracy, compared to full-graph training, regardless of the choice of the dataset, GNN models, and hardware settings. The study also shows that using time-to-accuracy as a metric is more relevant than comparing epoch times for systems that have different convergence criteria. Systems of both types can achieve similar accuracy when hyperparameter tuning is performed separately for full-graph and mini-batch training.
Our work stresses the importance of comparing GNN training systems of both classes in future work in this area. Using time-to-accuracy as the main metric for comparison across areas is important, as is performing separate hyperparameter tuning for the two training methods. These findings are relevant to both machine learning practitioners who utilize GNN training systems and systems researchers involved in their design and evaluation.
Appendix
The results for highest achievable accuracy for GAT and GCN model are shown in Table 11 and Table 12. Given a dataset, we must compare the accuracies obtained for two combinations: (FG-search, FG-train) for full-graph training and (MB-search, MB-train) for mini-batch training. The results are averaged after 5 runs.
Table 13 shows the best architectures obtained for every model-dataset combination.
Figure 7 and Figure 8 illustrate the best test accuracy when training the GCN and GAT respectively. The top row shows mini-batch training experiments and the bottom row shows synchronous full-graph training.
Table 11. Convergence for GAT model
Pubmed Ogbn-arxiv Reddit Ogbn-products
FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search
FG-train $77.70\pm 00.50$ $77.20\pm 00.34$ $71.74\pm 00.22$ $67.77\pm 00.31$ $95.17\pm 00.20$ $93.00\pm 00.23$ $74.43\pm 00.05$ $73.33\pm 00.15$
MB-train $75.80\pm 00.06$ $78.50\pm 00.03$ $70.19\pm 00.16$ $70.63\pm 00.12$ $94.58\pm 00.14$ $95.58\pm 00.13$ $74.27\pm 00.07$ $74.57\pm 00.07$
Table 12. Convergence for GCN model
Pubmed Ogbn-arxiv Reddit Ogbn-products
FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search
FG-train $79.40\pm 00.12$ $75.10\pm 00.37$ $71.54\pm 00.21$ $69.32\pm 00.22$ $94.67\pm 00.20$ $93.64\pm 00.10$ $75.71\pm 00.18$ $73.81\pm 00.13$
MB-train $78.20\pm 00.16$ $78.70\pm 00.23$ $70.54\pm 00.16$ $71.00\pm 00.16$ $93.31\pm 00.14$ $94.41\pm 00.14$ $75.27\pm 00.07$ $79.31\pm 00.04$
Table 13. Best architectures
Model Training technique Architecture Pubmed Ogbn-arxiv Reddit Ogbn-products
GraphSAGE FG Layers 4 3 4 5
Hidden size 64 512 1024 512
MB Layers 2 5 4 5
Hidden size 256 128 512 256
GAT FG Layers 3 3 2 2
Hidden size 1024 1024 1024 256
Num heads 1 2 2 2
MB Layers 2 4 2 3
Hidden size 1024 256 512 128
Num heads 4 2 2 2
GCN FG Layers 2 2 2 3
Hidden size 512 1024 1024 512
MB Layers 6 2 2 2
Hidden size 64 1024 512 512
Figure 7. Ablation study with number of layers and hidden size for GCN model. Legend represents the size of each hidden layer. X-axis shows the number of layers. Y-axis shows the test accuracy. Top row is for mini-batch training and bottom row is for full graph training. We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
Figure 8. Ablation study with number of layers and hidden size for GAT model. Legend represents the size of each hidden layer. X-axis shows the number of layers. Y-axis shows the test accuracy. Top row is for mini-batch training and bottom row is for full graph training. We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
Table 14 shows the results of the time to accuracy experiment for GAT model on a single host with NVLink connectivity.
Table 14. Time to accuracy for GAT on a single NVLink host with 4 GPUs. Architectures (#Layers, hidden size, #heads, fanout): ogbn-arxiv - (3, 1024, 6, 25), Reddit - (2, 1024, 4, 20), ogbn-products - (3, 128, 3, 20). ET - Epoch time, TTA - Time to accuracy. For ogbn-papers100m DGL uses UVA mode for sampling and Quiver performs distributed caching.
System ogbn-arxiv reddit ogbn-products
ET (s) PipeGCN (FG, Sync) 0.625 1.801 1.033
PipeGCN (FG, Async) 0.469 1.312 0.614
BNS-GCN (FG, Sync) 0.420 1.395 0.629
DGL (MB, Sync) 3.871 6.510 2.730
Quiver (MB, Sync) 3.596 6.779 2.974
TTA (s) PipeGCN (FG, Sync) 275.04 1728.58 413.28
PipeGCN (FG, Async) 211.14 1312.15 233.24
BNS-GCN (FG, Sync) 151.20 1395.40 132.13
DGL (MB, Sync) 42.58 97.65 40.95
Quiver (MB, Sync) 35.96 101.68 44.61
Table 15 shows the results for the scalability experiments for GAT model on PCIe setup and Table 16 shows the results of the scalability experiment on the NVLink setup.
Table 15. Single-host scalability with the GAT model using the PCIe host. Times in seconds. (#Layers, hidden size, #heads, fanout): ogbn-arxiv - (3, 1024, 2, 25), Reddit - (2, 1024, 1, 15), ogbn-products - (3, 128, 1, 10), Orkut - (2, 128, 1, 10). ET - Epoch time, TTA - Time To Accuracy.
Dataset ogbn-arxiv Reddit ogbn-products Orkut
System # of GPUs ET TTA TTA Speed-up ET TTA TTA Speed-up ET TTA TTA Speed-up ET
PipeGCN (Sync) 1 2.225 979.09 1.0x 5.857 5622.53 1.0x 3.510 1404.12 1.0x OOM
2 1.514 665.98 1.5x 4.035 3873.98 1.5x 3.381 1352.32 1.0x 2.726
3 1.323 581.98 1.7x 4.034 3872.83 1.5x 2.748 1099.16 1.3x 2.425
4 1.072 471.55 2.1x 2.890 2774.78 2.0x 2.452 980.60 1.4x 2.323
PipeGCN (Async) 1 2.218 998.01 1.0x 5.856 5856.10 1.0x 3.503 1331.10 1.0x OOM
2 1.334 600.08 1.7x 3.269 3269.20 1.8x 2.762 1049.37 1.3x 2.479
3 1.026 461.79 2.2x 3.109 3109.20 1.9x 1.972 749.17 1.8x 1.898
4 0.805 362.03 2.8x 1.879 1879.10 3.1x 1.523 578.78 2.3x 1.499
BNS-GCN (Sync) 1 2.337 841.43 1.0x 6.232 6231.60 1.0x 4.010 842.10 1.0x OOM
2 1.345 484.02 1.7x 3.472 3472.20 1.8x 3.466 727.82 1.2x 2.860
3 0.980 352.76 2.4x 2.697 2697.10 2.3x 2.366 496.92 1.7x 2.278
4 0.778 280.08 3.0x 2.193 2192.70 2.8x 1.710 359.06 2.3x 1.583
DGL (Sync) 1 15.401 154.01 1.0x 17.468 262.01 1.0x 18.394 367.88 1.0x 22.465
2 8.888 88.88 1.7x 8.575 128.62 2.0x 9.328 186.57 2.0x 11.580
3 6.146 61.46 2.5x 5.722 85.83 3.1x 6.246 131.17 2.8x 7.544
4 4.627 46.27 3.3x 4.594 68.91 3.8x 4.715 99.02 3.7x 6.444
Quiver (Sync) 1 16.230 162.30 1.0x 18.914 283.71 1.0x 17.457 349.15 1.0x 24.222
2 8.286 91.15 1.8x 9.002 135.03 2.1x 8.783 175.66 2.0x 14.134
3 5.737 68.85 2.4x 6.972 111.55 2.5x 5.633 118.29 3.0x 9.048
4 4.256 51.08 3.2x 4.595 87.31 3.2x 4.695 98.59 3.5x 6.695
Table 16. Single-host scalability with the GraphSAGE model using the NVLink host. Times in seconds. (#Layers, hidden size, fanout): ogbn-arxiv - (3, 1024, 25), Reddit - (2, 1024, 20), ogbn-products - (3, 128, 20), Orkut - (3, 128, 5). ET - Epoch time, TTA - Time to Accuracy.
Dataset ogbn-arxiv Reddit ogbn-products Orkut
System # of GPUs ET TTA TTA Speed-up ET TTA TTA Speed-up ET TTA TTA Speed-up ET
PipeGCN (Sync) 1 0.181 39.90 1.0x 0.433 121.16 1.0x 0.270 107.90 1.0x 0.466
2 0.182 40.04 1.0x 0.746 208.78 0.6x 0.398 159.28 0.7x 0.452
3 0.315 69.24 0.6x 0.735 205.68 0.6x 0.306 122.34 0.9x 1.312
4 0.211 46.51 0.9x 0.776 217.40 0.6x 0.518 207.20 0.5x 0.995
PipeGCN (Async) 1 0.182 56.28 1.0x 0.432 120.94 1.0x 0.270 102.48 1.0x 0.467
2 0.098 30.38 1.9x 0.255 71.41 1.7x 0.164 62.24 1.6x 0.281
3 0.071 21.97 2.6x 0.212 59.37 2.0x 0.114 43.17 2.4x 0.234
4 0.057 17.53 3.2x 0.130 36.42 3.3x 0.089 33.71 3.0x 0.183
BNS-GCN (Sync) 1 0.188 50.74 1.0x 0.369 58.97 1.0x 0.363 76.31 1.0x 0.536
2 0.106 28.71 1.8x 0.227 36.29 1.6x 0.219 46.03 1.7x 0.322
3 0.083 22.45 2.3x 0.183 29.33 2.0x 0.162 34.00 2.2x 0.289
4 0.069 18.56 2.7x 0.157 25.07 2.4x 0.136 28.64 2.7x 0.225
DGL (Sync) 1 1.519 7.60 1.0x 1.571 7.85 1.0x 2.819 42.28 1.0x 7.182
2 0.842 4.21 1.8x 0.846 4.23 1.9x 1.389 20.84 2.0x 3.606
3 0.577 2.89 2.6x 0.576 2.88 2.7x 0.926 13.90 3.0x 2.390
4 0.457 2.28 3.3x 0.490 2.45 3.2x 0.699 10.49 4.0x 1.842
Quiver (Sync) 1 0.906 9.06 1.0x 2.478 12.39 1.0x 2.463 36.94 1.0x 6.301
2 0.481 4.81 1.9x 1.278 6.39 1.9x 1.250 18.75 2.0x 3.317
3 0.333 3.33 2.7x 0.901 4.51 2.8x 0.861 12.92 2.9x 2.279
4 0.260 2.60 3.5x 0.731 3.66 3.4x 0.643 9.65 3.8x 1.743

### v3

@@
-Graph Neural Network Training Systems: A Performance Comparison of Full-Graph and Mini-Batch. [Experiment, Analysis & Benchmark]
+
 Abstract.
@@
 In this paper, we provide a comprehensive empirical comparison of representative full-graph and mini-batch GNN training systems.
-We find that the mini-batch training systems consistently converge faster than the full-graph training ones across multiple datasets, GNN models, and system configurations, with speedups between 2.4×2.4$\times$2.4 × - 15.2×15.2$\times$15.2 ×.
-We also find that both training techniques converge to similar accuracy values.
-Our work highlights the importance of comparing system across different classes, using time-to-accuracy rather than epoch time for performance comparison, and selecting appropriate hyperparameters for the comparison.
-PVLDB Reference Format:
-PVLDB, 14(1): XXX-XXX, 2025.
-doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
-doi:XX.XX/XXX.XX
-PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at https://github.com/saurabhbajaj123/GNN_minibatch_vs_fullbatch.
+We find that the mini-batch training systems consistently converge faster than the full-graph training ones across multiple datasets, GNN models, and system configurations.
+We also find that mini-batch training techniques converge to similar or often higher accuracy values as full-graph training ones, showing that mini-batch sampling is not necessarily detrimental to accuracy.
+Our work highlights the importance of comparing systems across different classes, using time-to-accuracy rather than epoch time for performance comparison, and selecting appropriate hyperparameters for each training method separately.
 1. Introduction
@@
 GNNs can be trained using either a mini-batch or a full-batch (typically called full-graph) approach, much like other machine learning models.
-In GNN training, however, these two approaches result in different message-passing patterns, which require different system pipelines and pose different system optimization challenges.
-This resulted in the development of two distinct classes of GNN training systems, each designed to support scalability for either mini-batch or full-graph training.
+In standard deep neural network (DNN) training, the dataset consists of individual training examples that can be processed independently and have no structural dependencies.
+In GNNs, in contrast, the training data is composed of vertices that are interconnected through edges, forming a graph structure where vertices cannot be treated as independent training examples.
+Full-graph and mini-batch training deal with these dependencies with different data management pipelines to partition data and parallelize computation and communication when scaling to large graphs.
+This resulted in the development of two distinct classes of GNN training systems, each designed to support either mini-batch or full-graph training.
 Figure 1. Different data management pipelines in two example systems: PipeGCN (full-graph) and DGL (mini-batch). The diagrams omit gradient synchronization.
-The main reason for the emergence of these two separate class of systems is that different training methods (full-graph vs.
-mini-batch) require different data management pipelines to partition data, transfer it across CPUs and GPUs, and parallelize computation when scaling to large graphs.
-Full-graph GNN training performs message-passing for the entire batch, typically across the entire graph, at each epoch.
-To scale to large graphs that do not fit in the memory of a single GPU, multi-GPU full-graph training systems use model parallelism: they partition the graph, process different partitions at different GPUs, and perform message passing across partitions (Tripathy et al., 2020; Ma et al., 2019; Jia et al., 2020).
-For example, the PipeGCN (Wan et al., 2022b) full-graph training system partitions the input graph and keeps each partition in a different GPU (see Figure 1a).
-The black dotted lines show the communication of vertex features across GPUs during message passing at each GNN layer in the forward and backward pass.
+Full-graph GNN training performs message-passing across the entire graph at each epoch.
+To scale to large graphs that do not fit in the memory of a single GPU, multi-GPU full-graph training systems use model parallelism: they partition the graph, process different partitions at different GPUs, and exchange hidden vertex features across partitions (Tripathy et al., 2020; Ma et al., 2019; Jia et al., 2020).
+For example, the PipeGCN (Wan et al., 2022b) full-graph training system partitions the input graph and keeps each partition in a different GPU (see Figure 1(a)).
+The black dotted lines show the communication of hidden vertex features across GPUs at each GNN layer in the forward and backward pass.
 The communication happens through either a fast cross-GPU bus such as NVLink, if available, or PCIe.
@@
 A unique challenge when applying data parallelism to GNNs is that preparing mini-batches requires expensive pre-processing at each iteration.
-At each iteration, data-parallel mini-batch training systems sample the k-hop neighborhood of the vertices in the mini-batch, load each sample into a different GPU and perform local message passing only within each sample, without passing message across GPUs.
-For example, Figure  1b shows a training iteration in the popular DGL system.
-First, each GPU obtains a sample of the k-hop neighborhood of a subset of the vertices in the mini-batch.
-Sampling is preferably executed using the GPU since it is a computationally expensive step in itself (Jangda et al., 2021).
+At each iteration, data-parallel mini-batch training systems sample a subset of the k-hop neighbors of the vertices in the mini-batch, load each sample into a different GPU, and perform local message passing only within each sample, without passing message across GPUs.
+For example, Figure  1(b) shows a training iteration in the popular DGL system.
+First, each GPU obtains a sample for a subset of the vertices in the mini-batch.
+Graph sampling is preferably executed using the GPU since it is a computationally expensive step in itself (Jangda et al., 2021).
 Next, the sample and the input features of its vertices are loaded to the GPUs.
@@
 Full-graph training systems must deal with expensive and irregular computation and high communication costs at each epoch.
-They propose sophisticated scheduling and partitioning algorithms to address these bottlenecks (Md et al., 2021; Peng et al., 2022; Wan et al., 2022b; Thorpe et al., 2021; Wan et al., 2023a).
-Work can be partitioned and scheduled offline, since the graph is static.
-Some full-graph training systems also propose techniques to reduce the volume of message data transferred across GPUs and hosts (Wan et al., 2023b; Peng et al., 2022; Wan et al., 2022a).
-In data-parallel mini-batch training systems, instead, computation at each iteration is much more lightweight and there is no cross-GPU message passing.
+In mini-batch training systems, instead, computation at each iteration is much more lightweight and there is no cross-GPU exchange of hidden features at each layer.
 However, pre-processing is a major bottleneck.
 Samples are typically large, so sampling them and loading them into GPUs is expensive.
-Research on scalable mini-batch systems has focused on addressing these bottlenecks by accelerating mini-batch sampling (Yang et al., 2022b; Jangda et al., 2021; Pandey et al., 2020; Wang et al., 2021a; Gong et al., 2023; Cai et al., 2023; Zheng et al., 2022a, b) and reducing the cost of loading the input features by caching them in the memory of the GPUs (Lin et al., 2020; Yang et al., 2022b; Song et al., 2023b; Cai et al., 2023; Tan et al., 2023; Yang et al., 2022a; Liu et al., 2023; Polisetty et al., 2023).
+The common rationale for using full-graph GNN training is to ensure that the training algorithm sees all the required dependent data in the graph, which is assumed to be the key to achieving high accuracy (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019).
+Mini-batch GNN training introduces additional variability in the gradient estimation, even compared to regular mini-batch DNN training.
+Specifically, in regular DNN training, the stochasticity in gradient estimation primarily arises from using mini-batches, which are random subsets of the training data, in each iteration.
+In mini-batch GNN training, the graph sampling used to form micro-batches is an additional source of stochasticity beyond the sampling of vertices in the mini-batch, because the training algorithm is not exposed to all the required dependent data in the graph.
 Motivation: What is the state of the art?
-To understand the state of the art in the field it is necessary to compare systems across different classes.
+Despite their different approaches and bottlenecks, full-graph and mini-batch training systems share the same goal: efficient and accurate training of Graph Neural Networks (GNNs).
+It is therefore important to establish which GNN training systems, be they full-graph or mini-batch systems, converge faster and to a higher accuracy under different scenarios.
+There is, however, no clear answer to this question in the literature.
 We performed an exhaustive review of the existing literature and found that prior works that propose systems in one class compare them with other systems in the same class, with limited or no performance evaluation of systems in the other class.
-Furthermore, several papers assert the superiority of the training approach they support in terms of the best accuracy that can be achieved, resulting in conflicting claims.
 We report a detailed discussion of the evidence in the literature in Section  2.
-Overall, the existing literature makes it difficult to get a clear picture of the state of the art of the field.
+Overall, the existing literature makes it difficult to get a clear picture of the state-of-the-art of the field.
 Performance evaluation.
-In this paper, we address this open issue by performing a thorough performance comparison of representative state-of-the-art GNN training systems across the two classes (Wan et al., 2022b, a; Wang et al., 2019; Zheng et al., 2020; Tan et al., 2023).
-We chose systems that have a stable, publicly-available, and open-source implementation, are representative of the most important types of systems optimizations in their class, and implement all the standard GNN models (GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GAT (Veličković et al., 2018)) in their repository.
-Our results show that the mini-batch training systems are consistently faster than the full-graph training counterparts in reaching the same target accuracy across all datasets, GNN model, and hardware deployments we considered.
+In this paper, we address this open issue by performing a thorough performance comparison of representative state-of-the-art GNN training systems across the two classes.
+For full-graph training, we consider PipeGCN (Wan et al., 2022b), BNS-GCN (Wan et al., 2022a), and AdaQP (Wan et al., [n.d.]b).
+For mini-batch training, we consider DGL (Wang et al., 2019), DistDGL (Zheng et al., 2020), and Quiver (Tan et al., 2023), and three sampling algorithms: Neighborhood Sampling (Hamilton et al., 2017), ClusterGCN (Chiang et al., 2019), and GraphSaint (Zeng et al., 2019).
+Our results show that the mini-batch training systems are consistently faster than the full-graph training counterparts in reaching the same target accuracy across all datasets, GNN models, and hardware deployments we considered.
 These results are consistent with the evidence in the literature.
-Several papers that propose full-graph training systems use epoch time as the main metric to compare with mini-batch training systems, showing that the latter have a larger epoch time  (Wang et al., 2022; Wan et al., 2023a; Md et al., 2021; Chen et al., 2023).
+Several papers that propose full-graph training systems use epoch time, which is the average time needed to execute one epoch, as the main metric to compare with mini-batch training systems, showing that the latter has a larger epoch time  (Wang et al., 2022; Wan et al., 2023b; Md et al., 2021; Chen et al., 2023).
 Our evaluation confirms these results.
-However, our results also highlight that considering the epoch time alone only gives a partial picture and that time-to-accuracy is a better metric for comparison.
+However, our results also highlight that time-to-accuracy is a better metric for comparison.
 Mini-batch training typically requires fewer epochs to converge because it updates the model parameters multiple times per epoch, once at each iteration, whereas full-graph training performs only one update per epoch.
@@
 Our empirical results show that this holds consistently across all systems, deployments, models, and datasets we consider.
-Across all the datasets we considered, the fastest mini-batch training system is between 2.4×2.4$\times$2.4 × to 15.2×15.2$\times$15.2 × (6.3×6.3$\times$6.3 × on average) faster than the fastest full-graph systems in terms of time-to-accuracy.
-In terms of scalability, full-graph training and mini-batch training systems observe similar relative speedups when they scale from a single GPU to multiple GPUs.
-However, the mini-batch systems we consider have a faster time-to-accuracy on a single GPU or host, so they remain faster in absolute terms in all configurations.
-We report our performance results in Section 4.
+We report our performance results in Section 4 and support and generalize our empirical observations with an analysis in Section 6.
 Performance vs.
 accuracy.
-The time-to-accuracy experiments discussed previously measure the time it takes different systems to reach the same target accuracy, which we select by performing hyperparameter tuning, using the same model and dataset.
-However, using a system with a longer time-to-accuracy can be justifiable if a model can converge to a higher accuracy.
+Using a system with a longer time-to-accuracy can be justifiable if a model can converge to a higher accuracy.
 This is an important factor when navigating the performance trade-offs of using different GNN training systems.
-In fact, some papers proposing full-graph GNN systems assert that full-graph training achieves higher accuracy (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others proposing mini-batch training systems claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
-Our results show that, with proper hyperparameter tuning, both full-graph and mini-batch training reach a similar accuracy across different datasets and models.
-Which method achieves the highest accuracy depends on the dataset and GNN model but the accuracy gaps are small.
-In our evaluation, we matched or exceeded all the accuracy results in the literature on GNN training systems we reviewed.
-We observe that given a dataset and GNN model, if we use the best hyperparameters for one training method with a system using the other training method, the latter achieves lower accuracy.
-This dependence on the choice of hyperparameters may explain the contradicting results found in the literature.
-Our results underline the importance of conducting separate hyperparameter tuning for the two training methods when comparing performance and accuracy across systems using different training methods.
-We also find that finding good hyperaparameters is easier with mini-batch training because it is less sensitive to variations in the model architecture.
-This has an impact on the overall training time, since model training often involves performing hyperparameter search.
-The harder it is to find good hyperparameters, the longer the overall training time.
-We report our accuracy experiments in Section 5.
+Our results show that, with proper hyperparameter tuning, mini-batch training can reach a similar accuracy as, if not higher accuracy than, full-graph training across different datasets and models.
+In our evaluation, we matched or exceeded all the test accuracy results for the GNN models we considered in the literature we reviewed.
+We observe that given a dataset and GNN model, hyperparameters that yield high accuracy for one training method do not work as well for the other.
+This may explain the contradicting results found in the literature and underline the importance of conducting separate hyperparameter tuning for the two training methods.
+These empirical findings also imply that filtering information during graph aggregation is not necessarily detrimental to accuracy, which is against the assumption made by works advocating full-graph training approaches.
+We discuss these results in Section 5.
 Lessons learned.
-Our results show that the mini-batch training systems we consider consistently achieve a better performance than the full-graph training ones and a similar accuracy.
+Our results show that the mini-batch training systems we consider consistently achieve better performance than the full-graph training ones and have similar accuracy.
 We stress, however, that the goal of this study is not to establish the inherent superiority of one class of systems over the other.
 Our study is an empirical evaluation and the field is continuously evolving.
-From a methodological perspective, our work highlights the importance of comparing GNN training systems across different classes.
-It also indicates a set of principles on how to perform this comparison: use time-to-accuracy as primary metric for comparison rather than epoch time, and perform separate hyperparameter tuning for mini-batch and full-graph training when selecting a target accuracy for performance comparison or comparing convergence accuracy.
-Prior works do not always follow these principles, and this may explain the apparent contradiction in the existing results.
-Our work offers guidance to ML practitioners who use GNN training systems and to systems researchers who design and evaluate them.
-An extended version of this paper, which includes additional experimental results that we had to omit due to space constraints, is available at (Bajaj et al., 2024).
-To ease reproducibility, our study uses publicly available datasets, model implementations, and training systems.
-We make all our scripts publicly available.
-[^1] https://github.com/saurabhbajaj123/GNN_minibatch_vs_fullbatch
+Mini-batch training also has limitations, such as a slightly higher variance in accuracy between training runs and a high variance across different sampling algorithms.
+Overall, our work highlights that existing algorithmic optimizations such as sampling or asynchrony can improve efficiency with no or minimal accuracy degradation, representing a promising research direction to achieve further performance and accuracy improvements.
+From a methodological perspective, our work highlights the importance of comparing GNN training systems across different classes and it also indicates a set of principles on how to perform this comparison.
 2. Background and Motivation
-This section provides necessary background on GNN and GNN training systems to follow the rest of the paper. It then discusses the existing evidence in the literature.
+This section provides the necessary background on GNN and GNN training systems to follow the rest of the paper. It then discusses the existing evidence in the literature.
 2.1. Graph Neural Networks
 Given a graph $G(V,E)$, where each node $v\in V$ is represented with a feature vector $h_{v}^{0}$ (input feature of vertex $v$) and the edges between any two vertices $u$ and $v$ is $e_{uv}\in E$.
-A GNN at $l^{th}$ layer performs the following computation to get the representation of any vector,
+A GNN at $l^{th}$ layer performs the following computation to get the hidden features of a vector:
 (1) $h_{v}^{l}=\Phi^{l}(h_{v}^{l-1},\Psi^{l}(h_{u}^{l-1},\forall u\in N\cup{v}),W^{l}),\$
-where $h_{v}^{l}$ is the $l^{th}$ layer representation for vertex $v$, $N$ represents the incoming neighbor vertices of vertex $v$, $W^{l}$ is the model weight matrix, $\Psi^{l}$ is any aggregation function and $\Phi^{l}$ is an update function.
-This paper considers the three standard GNN models used in previous work: GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2017).
-GraphSAGE learns the representation of a target node based on the features of the nodes in its k-hop neighborhood.
-The information from the neighboring nodes is propagated to the target node by performing message passing and aggregation recursively, one hop at a time.
-GAT (Graph Attention Networks) extend message passing by drawing inspiration from attention architectures in transformer models.
-It uses an attention score to weigh the incoming message of a node.
-GCN learns hidden layers representations and encode the graph structure and the features of the nodes by implementing a first-order approximation of spectral convolutions on graphs.
+where $h_{v}^{l}$ are the $l^{th}$ layer features for vertex $v$, $N$ represents the incoming neighbor vertices of vertex $v$, $W^{l}$ is the model weight matrix, $\Psi^{l}$ is any aggregation function and $\Phi^{l}$ is an update function.
 Table 1. Review of literature reporting claims of how mini-batch and full-graph training compare. (Notation: TTA - time to accuracy, ACC - accuracy, FG - Full graph, MB - mini-batch, ET - Epoch time)
-Proposes System/paper Main Claims Experimental evidence Year Open Source Platform
+Proposes System/paper Main FG vs. MB Claims Experimental evidence Year Open Source Evaluation Setup
 Full-graph GNNPipe (Chen et al., 2023) FG lower ET ET 2023 Multi-host multi-GPU
@@
 HongTu (Wang et al., 2023) MB information loss, MB lower ACC ACC, TTA 2023 Single-host multi-GPU, distributed CPU-only
-AdGNN (Song et al., 2023a) FG higher ACC, FG better convergence 2023 Multi-host single-GPU
-G3 (Wan et al., 2023a) FG higher ACC, lower ET ACC, ET 2023 Multi-host multi-GPU
-AdaQP (Wan et al., 2023b) MB information loss 2023 Mutli-host mult-GPU
-NeutronStar (Wang et al., 2022) FG lower ET ET 2022 Multi-host single-GPU
+ADGNN (Song et al., 2023a) FG higher ACC, FG better convergence 2023 Multi-host CPU, Multi-host single-GPU
+G3 (Wan et al., 2023b) FG higher ACC, lower ET ACC, ET 2023 Multi-host multi-GPU
+AdaQP (Wan et al., 2023c) MB information loss 2023 Multi-host multi-GPU
 PipeGCN (Wan et al., 2022b) FG higher ACC ACC 2022 Multi-host multi-GPU
 BNS-GCN (Wan et al., 2022a) FG higher ACC, lower ET ACC, ET 2022 Multi-host multi-GPU
+NeutronStar (Wang et al., 2022) FG lower ET ET 2022 Multi-host single-GPU
 Sancus (Peng et al., 2022) MB information loss 2022 Multi-host multi-GPU
@@
 DistGNN (Md et al., 2021) FG lower ET ET 2021 Multi-host CPU
-DGCL (Cai et al., 2021) MB ACC loss 2021 Mutli-host mult-GPU
+DGCL (Cai et al., 2021) MB ACC loss 2021 Multi-host multi-GPU
 CAGNET (Tripathy et al., 2020) FG higher ACC 2020 Multi-host multi-GPU
@@
 Full-graph and mini-batch training systems implement different data management pipelines for scalable multi-GPU training, which resulted in the emergence of two separate categories of systems.
-For a more detailed discussion on distributed GNN training systems, we refer the reader to (Shao et al., 2024; Lin et al., 2023).
-Full-graph training systems The key idea behind full-graph training systems is to partition a large graph into multiple subgraphs such that each can fit into a GPU, and then train them in parallel.
-This requires communication of vertex features across GPUs.
-Different types of communication strategies have been used in the literature, including broadcast-based (Tripathy et al., 2020), point-to-point based (Md et al., 2021; Demirci et al., 2022), pipeline-based (Wan et al., 2023a; Mostafa, 2022; Md et al., 2021; Wang et al., 2021b), and communication via shared memory (Jia et al., 2020; Ma et al., 2019).
-Communication incurs a substantial performance overhead and requires frequent synchronization.
-To overcome this bottleneck, two main types of optimizations have been proposed: asynchronous communication and sampling.
-Asynchronous communication allows the use of stale local information to avoid expensive synchronization costs.
-They can be categorized as epoch-fixed (Md et al., 2021; Wan et al., 2022b) and epoch-adaptive (Thorpe et al., 2021) staleness models.
-In GNN training, two synchronization points exist which are specific to GNN training: first, embedding aggregation and second, gradient aggregation.
-Some protocols (Peng et al., 2022; Md et al., 2021) only consider the embedding aggregation stage to support asynchronous aggregation, while others (Wan et al., 2022b; Thorpe et al., 2021) are designed to use historical information in both the embedding aggregation and gradient aggregation stages.
-Sampling keeps training synchronous but reduces communication cost by exchanging messages only among a subset of the boundary vertices across partitions (Wan et al., 2022a).
-Synchronous training is preferred in many cases because its results are predictable and reproducible.
-Mini-batch training systems Mini-batch training systems rely on the generation of mini-batches to scale GNNs to large graphs and employ graph sampling to reduce the neighborhood explosion problem.
-Mini-batch training systems have a different set of performance bottlenecks than full-graph training ones.
-Existing research has focused on optimizing two steps, mini-batch sampling and input feature loading, which are major performance bottlenecks and often have a cost comparable to training.
-Multiple vertex-wise, layer-wise and subgraph-wise sampling techniques have been proposed (Chen et al., 2018a; Hamilton et al., 2017; Zeng et al., 2019; Zou et al., 2019).
-It is common to accelerate these algorithms using GPUs (Jangda et al., 2021; Pandey et al., 2020; Wang et al., 2021a; Gong et al., 2023; Yang et al., 2022b).
-To avoid massive CPU-GPU communication while loading, some works (Zhu et al., 2019) have proposed caching techniques, like replicating frequently accessed remote vertices at local , store high out-degree vertices (Lin et al., 2020), caching hot vertices (i.e., those accessed more often) (Min et al., 2022).
-Some works (Liu et al., 2023) propose dynamic caching using a FIFO policy.
-Some distributed sampling techniques (Jiang and Rumi, 2021; Cai et al., 2023) prioritize local vertices, pushing sampling tasks to remote workers only if results need to be sent back, reducing communication.
-Some cache both topological and feature data on the GPUs (Sun et al., 2023; Zhang et al., 2023; Cai et al., 2023).
-To avoid graph sampling becoming a bottleneck, some works have focused on developing efficient GNN dataloaders (Bai et al., 2021; Deutsch et al., 2019; Dong et al., 2021; Min et al., 2021).
-Some systems reduce the cost of loading by proposing hybrid parallelism approaches for mini-batch training that are alternative to data parallelism: push-pull parallelism (Gandhi and Iyer, 2021) and split parallelism (Polisetty et al., 2023).
+In this paper, we focus our evaluation on GPU-based training on multi-GPU systems, which is a common configuration for scalable GNN training.
+For a comprehensive and detailed discussion on distributed GNN training systems, see also (Shao et al., 2024; Lin et al., 2023).
+Full-graph training systems.
+The key idea of full-graph training is to execute the GNN layers for all vertices in the graph, as shown in Eqn.
+1.
+To scale to large graphs, distributed full-graph training systems partition the graph into multiple subgraphs such that each can fit into the memory of one GPU.
+They use a model-parallel approach to train across GPUs that requires exchanging vertex features across partitions.
+The main focus of work on full-graph training systems has been on reducing the cost of vertex feature communication.
+We now discuss existing full-graph GNN training systems work in terms of two design dimensions: training data management and algorithmic optimizations (see Figure 2).
+Some full-graph training systems proposed optimizations that focused on the management of training data and do not introduce approximations that can potentially impact accuracy.
+If there are not enough GPUs to store each partition in a different GPU, it is necessary to transfer data between the host memory and GPU memory.
+Prior work such as NeuGraph (Ma et al., 2019), RoC (Jia et al., 2020), G3 (Wan et al., 2023b), and HongTu (Wang et al., 2023) have proposed techniques to optimize host-GPU communication.
+When there are enough GPUs to distribute all the partitions of the graph in GPU memory, training proceeds as depicted in Figure 1(a) (Wang et al., 2022; Wan et al., 2023b; Cai et al., 2021; Tripathy et al., 2020).
+Some recent works improve performance by further introducing algorithmic optimizations that can impact accuracy, unlike the work we described previously.
+Sancus (Peng et al., 2022), PipeGCN (Wan et al., 2022b), and GNNPipe (Chen et al., 2023) introduce asynchrony and allow GPUs to operate on stale vertex features.
+This enables GPUs to overlap communication with computation instead of having to wait at each layer for fresh vertex features coming from other GPUs.
+An alternative research direction has been to use sampling.
+BNS-GCN samples boundary nodes with edges across partitions at each epoch and exchanges vertex features only for those vertices (Wan et al., 2022a).
+ADGNN proposes an aggregation-difference sampling algorithm (Song et al., 2023a).
+Finally, the AdaQP system proposed message quantization to reduce the communication cost, together with new partitioning algorithms (Wan et al., 2023c).
+Figure 2. Classes of optimizations that impact performance (P) and accuracy (A), and representative systems evaluated in this work.
+Mini-batch training systems.
+Mini-batch training systems break up the training dataset into mini-batches and train and update the model on one mini-batch at a time.
+They use data parallelism to scale to multiple GPUs.
+Each GPU computes the hidden features of a subset of the vertices in the mini-batch, which is called the micro-batch.
+In principle, to execute $k$ GNN layers without loss of information as shown in Eqn.
+1, each GPU would need to load all the vertices in the $k$-hop neighborhood of the vertices in the micro-batch.
+The resulting subgraph, however, is often too large to be loaded into one GPU, inducing the so-called neighborhood explosion problem.
+Mini-batch training systems use sampling to load only a subset of the $k$-hop neighborhood.
+This sample-load-train pipeline is depicted in Figure 1(b).
+Compared to full-graph training, data parallelism eliminates the need to exchange hidden vertex features across partitions.
+The pipeline, however, introduces two different bottlenecks: sampling and data loading.
+We now discuss systems-level optimizations that mitigate these bottlenecks without impacting accuracy and then discuss the choice of sampling algorithms (see Fig 2).
+Initial mini-batch systems ran CPU-based sampling, but this can represent a major performance bottleneck, motivating the need for GPU-based sampling (Jangda et al., 2021; Serafini and Guan, 2021).
+Several systems-level optimizations have been proposed to run sampling algorithms on GPUs efficiently (Jangda et al., 2021; Pandey et al., 2020; Wang et al., 2021; Gong et al., 2023; Cai et al., 2023).
+In terms of management of the training data, loading the training data to the GPUs can be a significant performance bottleneck.
+To address it, prior work proposed caching training data in GPU memory, such as the input features (Lin et al., 2020) or the graph structure (Yang et al., 2022b; Zhang et al., 2023; Cai et al., 2023).
+In local GPU caching, each GPU only accesses its local cache (Lin et al., 2020; Yang et al., 2022b).
+In distributed GPU caching GPUs can directly access data cached in the main memory of other GPUs, which is faster than accessing data from the host memory when GPUs are connected through a fast bus like NVLink (Tan et al., 2023; Yang et al., 2022a; Song et al., 2023b; Cai et al., 2023).
+Mini-batch training systems can leverage GPU caching also when only a subset of the training dataset is cached (partial caching).
+They load data from the host memory in case of cache misses (Min et al., 2022; Yang et al., 2022b; Sun et al., 2023; Tan et al., 2023; Liu et al., 2023; Cai et al., 2023).
+In terms of algorithmic optimizations that can impact accuracy, mini-batch training systems can run different sampling algorithms.
+The seminal work that proposed mini-batch GNN training, GraphSage, proposed a simple neighborhood sampling approach (Hamilton et al., 2017).
+Subsequent work proposed many different algorithms such as FastGCN (Chen et al., 2018a), ClusterGCN (Chiang et al., 2019), LADIES (Zou et al., 2019), or GraphSaint (Zeng et al., 2019).
 2.3. Motivation
 The motivation for this work is that prior work on GNN training systems makes it difficult to establish the state of the art in the field.
-A significant amount of works on GNN training do not compare or evaluate their systems against systems using a different training approach(full-graph or mini-batch) (Kaler et al., 2023; Sun et al., 2023; Cai et al., 2023; Chen et al., 2018a; Hamilton et al., 2017; Chen et al., 2018b; Huang et al., 2018; Zeng et al., 2019; Gasteiger et al., 2022; Wang et al., 2021a; Zhu et al., 2019; Zhang et al., 2020; Kipf and Welling, 2017; Gandhi and Iyer, 2021).
-In the following, we review 33 papers on GNN training that do make claims comparing systems in the two classes.
+A significant amount of work on GNN training do not compare or evaluate their systems against systems using a different training approach and training pipeline (full-graph or mini-batch) (Kaler et al., 2023; Sun et al., 2023; Cai et al., 2023; Chen et al., 2018a; Hamilton et al., 2017; Chen et al., 2018b; Huang et al., 2018; Zeng et al., 2019; Gasteiger et al., 2022; Wang et al., 2021; Zhu et al., 2019; Zhang et al., 2020; Kipf and Welling, 2017; Gandhi and Iyer, 2021).
+Some prior work focuses only on the data-parallel mini-batching pipeline of Figure 1(b) and compares the performance of different low-level data management optimizations on top of the same underlying system (Yuan et al., 2024).
+In the following, we review 33 papers on GNN training that make claims comparing systems in the two classes.
 We summarize the claims in these papers in Table 1.
@@
 The last group consists of 4 papers that do not propose one specific training approach.
-Overall, we find the following shortcomings in the existing literature, which motivate our comprehensive experimental comparison.
+Overall, we find the following shortcomings in the existing literature, which motivate our work.
 Conflicting claims about how the two approaches compare.
-Several papers proposing full-graph systems (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wang et al., 2023; Song et al., 2023a) demonstrate that their systems achieve higher accuracy and lower epoch time compared to mini-batch systems.
-Some papers mention that mini-batch causes information loss, noisy gradients, and does not guarantee convergence (Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019; Wang et al., 2023; Song et al., 2023a) However, papers proposing mini-batch systems (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020) and some other works (Hu et al., 2020) present contradicting evidence and claims showing that mini-batch training leads to higher accuracy and faster convergence to the target accuracy level.
-Other works claim full-graph training is impractical, requires high memory and is hard to scale (Ai et al., 2023; Yang et al., 2022b; Zheng et al., 2022a; Chiang et al., 2019; Yao and Li, 2021).
-Some argue that no approach is clearly superior and assessing their effectiveness requires a comprehensive evaluation which is out of scope of their work (Liu et al., 2021; Kurt et al., 2023; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Liu et al., 2023; Lin et al., 2020; Wang et al., 2023).
+Several papers proposing full-graph systems (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wang et al., 2023; Song et al., 2023a) demonstrate that their systems achieve higher accuracy and lower epoch time compared to mini-batch systems.
+Some papers mention that mini-batch causes information loss, noisy gradients, and does not guarantee convergence (Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019; Wang et al., 2023; Song et al., 2023a) However, papers proposing mini-batch systems (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020) and some other works (Hu et al., 2020) present contradicting evidence and claims showing that mini-batch training leads to higher accuracy and faster convergence to the target accuracy level.
+Other works claim full-graph training is impractical, requires high memory, and is hard to scale (Ai et al., 2023; Yang et al., 2022b; Zheng et al., 2022a; Chiang et al., 2019; Yao and Li, 2021).
+Some argue that no approach is clearly superior and assessing their effectiveness requires a comprehensive evaluation which is out of the scope of their work (Liu et al., 2021; Kurt et al., 2023; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Liu et al., 2023; Lin et al., 2020; Wang et al., 2023).
 Comparing performance only in terms of epoch time.
-Certain works compare full-graph and mini-batch training systems based only on epoch time (Wang et al., 2022; Wan et al., 2023a; Md et al., 2021; Chen et al., 2023).
-As we will show, a comparison in terms of time to accuracy is more informative given the differences in the two training approaches.
+Certain works compare full-graph and mini-batch training systems based only on epoch time (Wang et al., 2022; Wan et al., 2023b; Md et al., 2021; Chen et al., 2023).
+As we will show, a comparison in terms of time-to-accuracy is more informative given the differences in the two training approaches.
 Our observations indicate that although mini-batch training exhibits higher epoch time, it still takes less time to reach a target accuracy level (Section 4.1).
@@
 This conflates the effect of the training approach with the GNN model choice, since different models may exhibit varying performance even when trained with the same technique.
-Some papers do not mention explicitly whether they perform separate hyperparameter tuning for each method (Wan et al., 2023a, 2022b, 2022a; Thorpe et al., 2021; Zhao et al., 2021; Kaler et al., 2022).
+Some papers do not mention explicitly whether they perform separate hyperparameter tuning for each method (Wan et al., 2023b, 2022b, 2022a; Thorpe et al., 2021; Zhao et al., 2021; Kaler et al., 2022).
 As we show in our experimental evaluation of Section 5, given a dataset and GNN model, it is possible to find hyperparameter settings where one method achieves higher accuracy than another and vice versa.
 Using hyperparameter tuning is necessary to perform a fair comparison.
-Table 2. Dataset statistics. (*) The input features and the training data for Orkut are synthetic
-Datasets #Nodes #Edges #Classes #Features Train/Val/Test
-Pubmed 19,717 108,365 3 500 18,217/500/1,000
-Ogbn-arxiv 169,343 1,166,243 40 128 90,941/29,799/48,603
-Reddit 232, 965 11, 606, 919 41 602 152,410/23,699/55,334
-Ogbn-products 2,449,029 61,859,140 47 100 195,922/48,980/2,204,127
-Orkut 3,072,441 117,185,083 2* 128* 1,843,464/614,488/614489*
-Ogbn-papers100M 111,059,956 1,615,685,872 172 128 1,207,179/125,265/214,338
-Table 3. Hyperparameter search space
-Parameter range/values
-Number of Hidden Layers [2, 10]
-Size of hidden layers [16, 1024]
-Aggregator (GraphSage only) { mean, gcn, pool }
-Number of heads (GAT only) [1, 12]
-Sampling Fanout (MB only) [4, 50]
-Learning rate [0.0001, 0.01]
-Dropout [0.2, 0.9]
-Mini-batch size (MB only) [128, 4096]
-3. Experimental setup
+Table 2. Dataset statistics. (*) The input features and the training data for Orkut are synthetic.
+Datasets Nodes Edges Classes Features Train/Val/Test
+Pubmed 19k 108k 3 500 18k/0.5k/1k
+Arxiv 169k 1.1M 40 128 91k/29k/48k
+Reddit 232k 11.6M 41 602 152k/23k/55k
+Products 2.4M 61.9M 47 100 196K/49K/2.2M
+Orkut 3M 117M 2* 128* 1.8M/614k/614k*
+Papers100M  [^1] M 1.6B 172 128 1.2M/125k/214k
+3. Experimental methodology
 Hardware setup.
@@
 By default, all our experiments use 4 GPUs per host.
-Datasets.
+Datasets and models.
 Our experiments use six datasets: pubmed (Namata et al., 2012), ogbn-arxiv (Hu et al., 2020), reddit (Hamilton et al., 2017), ogbn-products (Hu et al., 2020), orkut (Yang and Leskovec, 2012) and ogbn-papers100m (Hu et al., 2020) (see Table 2).
-Except orkut, all these datasets come with input features that enable evaluating the accuracy of GNN training.
+These datasets have varying average degrees, ranging from 5.6 to 60, and input feature sizes.
+They also differ on the fraction of vertices in the training set: in pubmed, this includes most vertices, whereas in papers it includes only 1.1% of the vertices.
+Except for orkut, all these datasets come with input features that enable evaluating the accuracy of GNN training.
 The orkut dataset contains only the graph topology, so we only use it to measure epoch time for the scalability experiments.
-To run GNN training on it, we add synthetic input features of size 128, a train/val/test split of 60%/20%/20%.
-One out of two target classes is randomly assigned to each node.
+We consider the three standard GNN models used in previous work: GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2017).
 Representative GNN training systems.
-To compare across full-graph and mini-batch distributed GNN systems, we select representative systems that incorporate the main optimizations discussed in Section 2.2.
-We chose systems that have a stable open-source implementation, including implementations of our target models (GCN, GraphSage, and GAT) in their repositories [^2] DGL 1.1.2: https://github.com/dmlc/dgl; Quiver 0.1.0: https://github.com/quiver-team/torch-quiver; DistDGL: https://github.com/dmlc/dgl/tree/master/examples/distributed; PipeGCN: https://github.com/GATECH-EIC/PipeGCN; BNS-GCN: https://github.com/GATECH-EIC/BNS-GCN.
+To compare across full-graph and mini-batch distributed GNN systems, we select representative systems that have publicly available and stable implementations and incorporate the main design choices discussed in Section 2.2, as summarized in Figure 2.
 We run all systems on top of PyTorch 2.0.1 and Python 3.8.10.
-For full-graph training, we select PipeGCN (Wan et al., 2022b), which supports asynchronous training to overlap communication and computation, and BNS-GCN (Wan et al., 2022a), which uses sampling to reduce communication.
-Both systems partition the entire graph topology and features, and store each partition in the memory of one GPU.
-For mini-batch training, we choose popular data-parallel systems that support GPU-based sampling, efficient input feature loading, and feature caching.
-Specifically, we use DGL (Wang et al., 2019), which only supports single-GPU caching, and Quiver (Tan et al., 2023), which supports distributed caching.
-DistDGL (Zheng et al., 2020) is built on top of DGL and performs mini-batch training on a distributed cluster of hosts.
-It only supports CPU-based training and does not support caching.
-All the mini-batch systems we consider support only synchronous training.
-Model selection and implementations.
-We consider standard GNN models that are commonly used to evaluate training systems: GCN, GraphSage, and GAT.
-We use the implementations provided with the distributions of the systems we benchmark and validate that the semantics of the implementations are consistent across different systems.
-This is to make sure that the model implementations are fully compliant and optimized for the respective systems and avoid introducing biases.
-DGL and Quiver include implementations of all the three models in its distribution.
-The PipeGCN distribution includes GraphSAGE, while the GCN and GAT implementations for both PipeGCN and BNS-GCN are included in the BNS-GCN distribution.
-For mini-batch training, we considered the standard Neighborhood Sampling algorithm for our evaluation (Hamilton et al., 2017).
-Hyperparameter search space.
-The search space for hyperparameters is shown in Table 3.
-The first set of hyperparameters define the architecture of the GNN models.
-Fanout is related to the neighborhood sampling algorithm used in mini-batch training.
-Finally, we consider general model training parameters such as learning rate, dropout and batch size.
-To search this parameter space, we use random search for all the parameters to obtain the best set of hyperparameters giving the highest validation accuracy.
+For full-graph training, in the training data management dimension of Figure 2, our evaluation only considers the more favorable situation where the entire training data is fully cached in the memory of the GPUs since they can entirely avoid the overhead of GPU-host communication.
+All systems adopt a pipeline similar to the one of Figure 1(a).
+The Full-Graph baseline is synchronous and does not use any optimization that can impact accuracy (Wan et al., 2023a).
+In the dimension of algorithmic optimization, we consider three baselines representing three classes of algorithmic optimizations: PipeGCN for asynchronous training (Wan et al., 2022b, 2023a), BNS-GCN for sampling (Wan et al., 2022a, [n.d.]a), and AdaQP for message quantization (Wan et al., 2023c, [n.d.]b).
+For mini-batch training, we consider three baselines: DGL (Wang et al., 2019; dgl, 2023), its distributed version DistDGL (Zheng et al., 2020; dgl, 2023), and Quiver (Tan et al., 2023; Quiver, 2023).
+These cover the system-level optimizations discussed in Section 2.2 (see Figure 2).
+All systems adopt a pipeline similar to the one of Figure 1(b).
+In terms of sampling implementations, DistDGL uses CPU-based sampling, while both DGL and Quiver support GPU-based sampling.
+DistDGL partitions the training dataset across multiple hosts.
+In terms of GPU caching, DGL only supports local GPU caching while Quiver also supports distributed GPU caching.
+Quiver uses distributed input feature caching for the Orkut and Papers100M datasets, which are too large for local caching.
+Mini-batch training can use different sampling algorithms to select a subset of the k-hop neighbors of the vertices in the mini-batch.
+We consider the standard Neighborhood Sampling (NS) algorithm, which is available for both DGL and Quiver, ClusterGCN (Chiang et al., 2019), and GraphSaint (Zeng et al., 2019).
+Hyperparameter search.
+We run an extensive hyperparameter tuning.
+We validated our search by matching the best test accuracies we found in the literature on GNN training systems we reviewed.
+To manage the large search space, we iterate multiple times over three phases, where we tune a set of hyperparameters and fix all others until we don’t observe any more improvements in accuracy.
+In the first phase, we use grid search for the architectural hyperparameters, namely the number of layers or the hidden feature size, which have a limited number of discrete values.
+In the second phase, we run a random search for the remaining hyperparameters, such as the learning rate.
+Mini-batch training requires tuning the micro-batch size and additional sampling hyperparameters compared to full-graph training.
+We tune these in a third phase using Bayesian optimization, reducing the search space by considering only a few commonly used discrete values.
+In the appendix, we report our hyperparameter space in table 11 and show that mini-batch training is less sensitive than full-graph training to the tuning of the architectural hyperparameters.
 4. Performance Evaluation
 We evaluate the performance of our representative GNN training systems and answer two questions: Q1) How long do representative systems take to reach the same target accuracy?
-Q2) How do these systems scale with a varying number of GPUs and hosts?
-We mainly use time-to-accuracy as the main metric for performance comparison since it is a more informative metric than epoch time, as we will discuss.
-We conduct end-to-end performance measurements because mini-batch and full-graph training systems have different implementations and use different optimizations.
-For a more in-depth discussion of the effect of specific systems optimizations, it is necessary to restrict the evaluation to one class of systems, as done in (Yuan et al., 2024) for mini-batch systems.
+(Section 4.1), and Q2) How do these systems scale with a varying number of GPUs and hosts?
+(Section 4.2).
 4.1. Time-to-Accuracy
-We start by answering question Q1: How long do representative systems take to reach the same target accuracy?
 Methodology Time-to-accuracy is the time taken by any system to train to a target accuracy.
@@
 We then use the same model architecture across all GNN training systems and measure their time to reach the target accuracy.
+The time-to-accuracy considers only the final training phase and not the hyperparameter search.
+For mini-batch systems, we consider the common Neighborhood Sampling algorithm as default because it is available on all baselines and it consistently achieves an accuracy close to the full-graph training methods.
+We report the hyperparameters in the appendix Appendix.
 These experiments consider the two host types described in Section 3: a higher-end NVLink host and a lower-end PCIe host.
-4.1.1. Single NVLink host
-Table 4.
-Time to accuracy for GraphSAGE on a single NVLink host with 4 GPUs.
-Architectures (# layers, hidden size): pubmed - (3,256), ogbn-arxiv - (2,512), reddit - (3,1024), ogbn-products - (3,256), ogbn-papers100m - (2, 128).
-ET - Epoch time, TTA - Time to accuracy.
-For ogbn-papers100m DGL uses UVA mode for sampling and Quiver performs distributed caching.
-(Best results shown in bold.)
-System pubmed ogbn-arxiv reddit ogbn-products ogbn-papers100m
-Epoch time (s) PipeGCN (FG, Sync) 0.009 0.057 1.845 2.086 1.094
-PipeGCN (FG, Async) 0.008 0.011 0.370 0.341 0.644
-BNS-GCN (FG, Sync) 0.012 0.017 0.483 0.473 0.846
-DGL (MB, Sync) 0.007 0.110 2.262 1.335 3.428
-Quiver (MB, Sync) 0.005 0.131 1.994 2.117 1.878
-Time to accuracy (s) PipeGCN (FG, Sync) 0.87 11.03 202.95 479.81 984.33
-PipeGCN (FG, Async) 1.02 3.44 103.47 122.64 553.75
-BNS-GCN (FG, Sync) 1.59 4.51 144.98 217.66 262.11
-DGL (MB, Sync) 0.44 1.10 6.78 46.73 75.42
-Quiver (MB, Sync) 0.36 0.65 9.97 84.70 41.31
-Table 5.
-Time to accuracy with 3 NVLink hosts using the ogbn-papers100m graph and a GraphSAGE model with 2 layers and a hidden size of 128 (DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached).
-(Best results shown in bold.)
-System Time (s)
-Epoch time (s) PipeGCN (FG, Sync) 1.139
-PipeGCN (FG, Async) 0.352
-BNS-GCN (FG, Sync) 0.432
-DistDGL (MB, Sync) 4.452
-Time to accuracy (s) PipeGCN (FG, Sync) 1093.63
-PipeGCN (FG, Async) 306.59
-BNS-GCN (FG, Sync) 133.92
-DistDGL (MB, Sync) 35.62
-Table 6.
-Time to accuracy and epoch time for GraphSAGE with a single PCIe host.
-Notation: ET - Epoch time, TTA - Time to accuracy.
-Architectures (# layers, hidden size): pubmed (3,256), ogbn-arxiv (2,512), reddit (4,1024) & ogbn-products (5,256).
-(Best results shown in bold.)
-System pubmed ogbn-arxiv reddit ogbn-products
-ET (s) PipeGCN (FG, Sync) 0.024 0.188 5.602 5.835
-PipeGCN (FG, Async) 0.020 0.067 3.278 2.526
-BNS-GCN (FG, Sync) 0.031 0.072 3.399 2.915
-DGL (MB, Sync) 0.019 0.314 9.416 7.081
-Quiver (MB, Sync) 0.019 0.216 10.272 9.246
-TTA (s) PipeGCN (FG, Sync) 2.40 36.16 616.25 1341.95
-PipeGCN (FG, Async) 2.35 20.75 917.80 909.43
-BNS-GCN (FG, Sync) 2.91 19.33 1019.64 1340.80
-DGL (MB, Sync) 1.25 3.14 28.25 247.83
-Quiver (MB, Sync) 1.32 1.08 51.36 369.85
-Table 7.
-Time to accuracy and epoch time with 3 PCIe hosts using the ogbn-papers100m graph and a GraphSAGE model with 2 layers and a hidden size of 128 (DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached).
-(Best results shown in bold.)
-System Time (s)
-Epoch time (s) PipeGCN (FG, Sync) 22.00
-PipeGCN (FG, Async) 1.03
-BNS-GCN (FG, Sync) 2.46
-DistDGL (MB, Sync) 38.37
-Time to accuracy (s) PipeGCN (FG, Sync) 19141.48
-PipeGCN (FG, Async) 956.34
-BNS-GCN (FG, Sync) 859.26
-DistDGL (MB, Sync) 767.38
-Figure 2. Time to accuracy for ogbn-arxiv, reddit, ogbn-products for GraphSAGE (single NVLink host).
-Figure 3. Time to accuracy for ogbn-papers100M on (3 NVLink hosts)
-Table 4 shows the time to accuracy when training GraphSAGE on a single NVLink host.
+Table 3. Time to accuracy (TTA) and epoch time (ET) for GraphSAGE (NVLink host).
+System pubmed ogbn- reddit ogbn- ogbn-
+arxiv products papers100m
+ET (s) Full-Graph 0.009 0.057 1.845 2.086 1.094
+PipeGCN 0.008 0.011 0.370 0.341 0.644
+BNS-GCN 0.012 0.017 0.483 0.473 0.846
+AdaQP 0.027 0.029 0.830 0.312 0.471
+DGL 0.007 0.110 2.262 1.335 3.428
+Quiver 0.005 0.131 1.994 2.117 1.878
+TTA (s) Full-Graph 0.87 11.03 202.95 479.81 984.33
+PipeGCN 1.02 3.44 103.47 122.64 553.75
+BNS-GCN 1.59 4.51 144.98 217.66 262.11
+AdaQP 0.81 19.86 41.5 53.04 160.41
+DGL 0.44 0.55 6.78 40.05 75.42
+Quiver 0.33 0.65 5.98 63.52 41.31
+Figure 3. Convergence curve for GraphSAGE (NVLink host).
+NVLink host.
+Table 3 and Figure 3 show the time to accuracy when training GraphSAGE on a single NVLink host with 4 GPUs.
 Full-graph training systems have a consistently lower epoch time (ET) than mini-batch training systems because they perform only one training iteration per epoch.
 However, their time-to-accuracy (TTA) is substantially higher compared to mini-batch systems.
-For example, on the largest graph we consider, ogbn-papers100m, Quiver is 6.4$\times$ faster than BNS-GCN.
-Similarly, for pubmed, Quiver is 2.4×2.4$\times$2.4 × faster than PipeGCN sync (lowest speed-up we observed) and for reddit, DGL is 15.2×15.2$\times$15.2 × faster than PipeGCN async (highest speed-up we observed).
-On average, the fastest mini-batch systems are 6.3×6.3$\times$6.3 × faster than the fastest full-graph system we considered for these datasets.
-This is because mini-batch training systems perform multiple training iterations and model updates per epoch and thus converge in fewer epochs.
-Figure 2 shows the time-to-accuracy curves.
-The dotted line indicates the target accuracy.
-We observe mini-batch training consistently has a much faster time to accuracy across graphs with different sizes.
-Results on other models are shown in the extended version (Bajaj et al., 2024).
-The experiments also show the impact of the systems optimizations introduced in previous work.
-Among the full-graph training systems, PipeGCN can substantially decrease epoch time and time-to-accuracy using asynchronous training compared to its baseline synchronous version.
-Sampling, as proposed by BNS-GCN, is less effective than asynchronous training on the smaller graphs, but it yields the fastest time-to-accuracy among all full-graph training systems on ogbn-papers100m.
-All full-graph systems we consider fully cache the graph in GPU memory.
-Mini-batch training systems also cache all the graph in the GPU memory with the exception of ogbn-papers100m, which cannot be cached by DGL.
-This is because DGL only supports caching when the entire graph fits in the memory of each GPU, unlike Quiver which implements a distributed cache where the graph can be partitioned among multiple GPUs.
-Despite the lack of caching, DGL still achieves a lower time-to-accuracy than the full-graph training systems on ogbn-paper100m.
-4.1.2. Multiple NVLink hosts
-We perform a similar experiment using three NVLink hosts with the largest of the graphs we consider, ogbn-papers100m.
-Table 5 shows the results.
-Figure 3 shows the corresponding time-to-accuracy curves.
-Quiver does not support multi-node training, so we do not include it in these experiments.
-We also replace DGL with its distributed version, DistDGL.
-The main trends are similar to the ones observed for the single-host experiments.
-DistDGL, which uses mini-batch training, has a much larger epoch time than the other systems, which use full-batch training.
-Its time-to-accuracy, however, is much shorter than the other systems.
-It is worth noting that DistDGL has a much larger time-to-accuracy than DGL on a single host (see Table 4).
-This is because DistDGL uses CPU-based sampling and does not support caching, unlike DGL.
-4.1.3. PCIe hosts
-Table 6 reports the results for single-host training with GraphSAGE using a lower-end host setup, where NVLink is not available and cross-GPU communication within one host must occur over a slower PCIe bus.
-Compared to the NVLink experiments of Table 4, the epoch time of all systems is larger because the GPUs are less powerful.
-Full-graph training systems experience a stronger degradation in performance than DGL because they must run cross-GPU communication over PCIe, whereas DGL avoids communication by using data parallelism.
-Quiver also observes a performance degradation since loading data from other GPU’s cache is not as effective over the PCIe bus as it is over NVLink.
-Table 7 futher compares all systems using 3 PCIe hosts for the ogbn-papers100m graph.
-In the PCIe setup, full-graph training systems need a multi-node setup to fully cache the ogbn-papers100m graph, which is larger than the others.
-In the setup with 3 PCIe hosts, hosts are connected by a slower network (1 GBps instead of 25 GBps).
-Like in the previous experiments, DistDGL has a lower time-to-accuracy than the other full-graph training systems, but the difference is much smaller than with the NVLink setup.
-This is because DistDGL requires cross-host communication to perform sampling and feature loading.
-These costs become bottlenecks over the slow 1GBps network of the PCIe host.
-By contrast, full-graph training systems are optimized to minimize communication and amortize its cost.
+For example, on the largest graph we consider, ogbn-papers100m, Quiver is 3.9×3.9$\times$3.9 × faster than AdaQP.
+Similarly, for pubmed, Quiver is 2.5×2.5$\times$2.5 × faster than AdaQP; for arxiv, DGL is 6.3×6.3$\times$6.3 × faster than PipeGCN; for reddit, Quiver is 6.9×6.9$\times$6.9 × faster than AdaQP (highest speed-up we observed); for products, DGL is 33% faster than AdaQP (lowest speed-up we observed).
+Results on other models are shown in the appendix in table 17.
+Comparing the performance of systems in the same class shows the impact of the systems optimizations introduced in previous work (see Figure 2).
+All full-graph systems we consider benefit from fully caching the graph in GPU memory.
+PipeGCN can substantially decrease ET and TTA compared to Full-Graph.
+By using asynchronous training and overlapping communication with computation, PipeGCN can mitigate the cost of synchronous cross-GPU vertex feature communication at each layer, which is high even when NVLink.
+Boundary-node sampling, as proposed by BNS-GCN, is less effective than asynchronous training on the smaller graphs, but it yields the fastest TTA among all full-graph training systems on ogbn-papers100M.
+The use of of synchronous communication in BNS-GCN results in larger ET than PipeGCN for all datasets, but using fresh training data contributes to achieving a shorter number of epochs to convergence with papers100M.
+AdaQP also combines synchrony with reduced communication costs thanks to message quantization.
+It speeds up ET significantly compared to Full-Graph, but not compared to other full-graph optimizations.
+Nonetheless, AdaQP can have faster TTA than other full-graph graph systems, since for some datasets it requires fewer epochs to converge.
+Among mini-batch systems, DGL supports local full caching and can cache all the graphs in Table 3 except papers100m.
+When full local caching is possible, DGL does not need to transfer input vertex features to the GPUs, maximizing GPU utilization.
+Quiver also uses full local caching, achieving a similar ET and TTA.
+However, Quiver implements a partial distributed cache across multiple GPUs that can also cache ogbn-papers100m.
+With distributed caching, some input vertex features may still be transferred among GPUs using NVLink buses, which is cheaper than loading the features from the host memory over the slower PCIe bus.
+This is why Quiver has a much lower time-to-accuracy than DGL for papers100M.
+Table 4. Time to accuracy (TTA) and epoch time (ET) for GraphSAGE (PCIe host).
+System pubmed ogbn-arxiv reddit ogbn-products
+ET (s) Full-Graph 0.024 0.188 5.602 5.835
+PipeGCN 0.020 0.067 3.278 2.526
+BNS-GCN 0.031 0.072 3.399 2.915
+DGL 0.019 0.314 9.416 7.081
+Quiver 0.019 0.216 10.272 9.246
+TTA (s) Full-Graph 2.40 36.16 616.25 1341.95
+PipeGCN 2.35 20.75 917.80 909.43
+BNS-GCN 2.91 19.33 1019.64 1340.80
+DGL 1.25 1.57 28.25 247.83
+Quiver 1.22 1.08 30.82 323.61
+PCIe host.
+Table 4 reports the results for single-host training with GraphSAGE using a lower-end host setup, where NVLink is not available and all cross-GPU communication must occur over a slower PCIe bus.
+The 4 GPUs are computationally less powerful and have less memory.
+The papers100M graph is too large to be fully cached in a single host by our full-graph training systems.
+We cannot run AdaQP since its implementation is not compatible with our lower-end GPUs.
+Like in the previous hardware setups, mini-batch training systems show a larger ET (with the exception of pubmed) but a shorter TTA on all datasets.
+Full-graph training systems must run cross-GPU vertex feature exchange over the slow PCIe bus.
+Optimizations that reduce the communication cost, like asynchrony (PipeGCN) and sampling (BNS-GCN), can reduce TTA, but the cost remains high.
+DGL and Quiver avoid vertex feature communication by replicating the training data on all GPUs so the advantage over the full-graph training systems is larger than with the NVLink host, reaching 32×32$\times$32 × for Reddit.
+The impact of dataset characteristics.
+Different datasets have an impact on the performance of different systems.
+While the impact on TTA is hard to understand due to the stochastic nature of training, the impact on ET can be more easily analyzed.
+The communication cost of full-graph training systems is determined by the number of boundary vertices across partitions.
+Reddit and products, for example, have a large number of boundary vertices, so the ET of the Full-Graph baseline is much larger compared to DGL and Quiver.
+Optimizations that reduce this communication cost used by PipeGCN and BNS-GCN can reduce this gap significantly.
+For mini-batch training, the training cost is determined by the size of the micro-batches assigned to each GPU.
+Datasets that require a larger number of GNN layers, such as products and reddit which requires 5 and 4 layers respectively, result in a larger ET for mini-batch training systems compared to the other full-graph training systems.
+In contrast, papers100M only requires 2 layers so the gap in ET is smaller.
+Different mini-batch sampling algorithms.
+Finally, we combine DGL with different mini-batch sampling algorithms than Neighborhood Sampling (NS).ClusterGCN achieves on average 2.8×2.8$\times$2.8 × faster TTA than NS across datasets, ranging from 0.14×0.14$\times$0.14 × slower to 12.3×12.3$\times$12.3 × faster depending on the dataset, and its ET is 6.4×6.4$\times$6.4 × faster on average.
+GraphSaint has 3.2×3.2$\times$3.2 × faster TTA than NS on average, ranging from 0.5×0.5$\times$0.5 × slower to 16.1×16.1$\times$16.1 ×, and faster ET by 9.3×9.3$\times$9.3 × on average (see appendix tables 24 and 25).
 Takeaway.
-Mini-batch training enables faster time-to-accuracy compared to full-graph training since it requires fewer epochs to reach a target accuracy.
-This is because full-graph training systems update the model only once per epoch, whereas mini-batch training systems perform multiple iterations per epoch and update the model at each iteration.
-By updating the model more frequently, mini-batch training systems typically converge faster than full-graph training ones, even if their epoch time is longer.
-These findings stress that time-to-accuracy measurements should always be included to provide a comprehensive comparison of systems using different training approaches.
-Comparing only in terms of epoch time, as is commonly done in prior work (see Table 1) does not provide a comprehensive comparison.
+Mini-batch training systems converge faster than full-graph training ones because they require fewer epochs to reach a target accuracy, despite having longer epochs.
+We observe this across all models, datasets, and hardware configurations we consider.
+Avoiding data transfers by caching input data in GPU memory has a strong impact on the performance of mini-batch training.
+For full-graph training, asynchrony (PipeGCN) and sampling (BNS-GCN) can reduce communication cost and thus ET, resulting in faster TTA.
+AdaQP can achieve even faster TTA even if its ET is larger than other full-graph methods, showing the importance of using algorithmic optimizations besides system-level optimizations.
+For mini-batch training, using different sampling algorithms can speed up TTA significantly in many cases.
 4.2. Scalability
-Table 8.
-Single-host scalability with the GAT model using the NVLink host.
-Times in seconds.
-(#Layers, hidden size, #heads, fanout): ogbn-arxiv - (3, 1024, 6, 25), Reddit - (2, 1024, 4, 20), ogbn-products - (3, 128, 3, 20), Orkut - (3, 128, 2, 5).
-ET - Epoch time, TTA - Time to Accuracy.
+Table 5. Single-host scalability for GAT (NVLink host). ET - Epoch time (s), TTA - Time to Accuracy (s).
 Dataset ogbn-arxiv reddit ogbn-products orkut
 System # of GPUs ET TTA TTA Speed-up ET TTA TTA Speed-up ET TTA TTA Speed-up ET ET speed-up
-PipeGCN (Sync) 1 1.293 569.01 1.0x 4.265 4094.59 1.0x 2.058 823.36 1.0x 1.898 1.0x
+Full-Graph 1 1.293 569.01 1.0x 4.265 4094.59 1.0x 2.058 823.36 1.0x 1.898 1.0x
 2 0.868 381.74 1.5x 2.566 2463.65 1.7x 1.341 536.20 1.5x 1.873 1.0x
@@
 4 0.625 275.04 2.1x 1.801 1728.58 2.4x 1.033 413.28 2.0x 1.442 1.3x
-PipeGCN (Async) 1 1.292 581.27 1.0x 4.268 4267.70 1.0x 2.060 782.76 1.0x 1.898 1.0x
+PipeGCN 1 1.292 581.27 1.0x 4.268 4267.70 1.0x 2.060 782.76 1.0x 1.898 1.0x
 2 0.777 349.83 1.7x 2.309 2309.20 1.8x 1.135 431.34 1.8x 1.149 1.7x
@@
 4 0.469 211.14 2.8x 1.312 1312.15 3.3x 0.614 233.24 3.4x 0.720 2.6x
-BNS-GCN (Sync) 1 1.309 471.06 1.0x 4.189 4188.60 1.0x 2.171 455.95 1.0x 2.173 1.0x
+BNS-GCN 1 1.309 471.06 1.0x 4.189 4188.60 1.0x 2.171 455.95 1.0x 2.173 1.0x
 2 0.764 274.93 1.7x 2.315 2315.30 1.8x 1.213 254.73 1.8x 1.284 1.7x
@@
 4 0.420 151.20 3.1x 1.395 1395.40 3.0x 0.629 132.13 3.5x 0.713 3.0x
-DGL (Sync) 1 15.006 150.06 1.0x 24.006 360.09 1.0x 10.554 158.31 1.0x 14.791 1.0x
+DGL 1 15.006 150.06 1.0x 24.006 360.09 1.0x 10.554 158.31 1.0x 14.791 1.0x
 2 7.290 72.90 2.1x 12.794 191.91 1.9x 5.827 87.40 1.8x 7.668 1.9x
@@
 4 3.871 42.58 3.5x 6.510 97.65 3.7x 2.730 40.95 3.9x 3.980 3.7x
-Quiver (Sync) 1 13.813 138.13 1.0x 26.427 396.41 1.0x 11.212 168.19 1.0x 13.718 1.0x
+Quiver 1 13.813 138.13 1.0x 26.427 396.41 1.0x 11.212 168.19 1.0x 13.718 1.0x
 2 6.575 65.75 2.1x 13.764 206.46 1.9x 5.797 86.95 1.9x 7.447 1.8x
@@
 4 3.596 35.96 3.8x 6.779 101.68 3.9x 2.974 44.61 3.8x 3.855 3.6x
-Table 9.
-Distributed scalability with the GraphSage model using 4 GPUs per host connected with NVLink .
-Times in seconds.
-(#Layers, hidden size, fanout): Orkut - (3, 512, 10), ogbn-products100M - (2, 512, 10).
-ET - Epoch time, TTA - Time to Accuracy.
-DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached
+Table 6. Distributed scalability for GraphSage (NVLink hosts). ET - Epoch time (s), TTA - Time to Accuracy (s).
 orkut ogbn-papers100m
 System # Hosts ET ET Speed-up ET TTA TTA Speed-up
-PipeGCN 1 4.769 1.0x 3.556 924.46 1.0x
-(Sync) 2 4.588 1.0x 3.566 927.13 1.0x
+Full-Graph 1 4.769 1.0x 3.556 924.46 1.0x
+2 4.588 1.0x 3.566 927.13 1.0x
 3 4.351 1.0x 3.933 1022.63 0.9x
 PipeGCN 1 0.454 1.0x 0.759 197.57 1.0x
-(Async) 2 0.341 1.3x 0.464 120.67 1.6x
+2 0.341 1.3x 0.464 120.67 1.6x
 3 0.227 2.0x 0.361 93.96 2.1x
 BNS-GCN 1 0.810 1.0x 0.997 209.38 1.0x
-(Sync) 2 0.608 1.3x 0.762 160.10 1.3x
+2 0.608 1.3x 0.762 160.10 1.3x
 3 0.405 2.0x 0.542 113.88 1.8x
-DistDGL 1 9.826 1.0x 11.561 80.93 1.0x
-(Sync) 2 7.305 1.3x 10.749 75.24 1.1x
-3 6.815 2.0x 6.612 46.28 1.7x
-Next we answer Q2: How do systems scale with a varying number of GPUs and a varying number of hosts?
-For each dataset, we pick a GNN architecture that fits in a single GPU and use the single-GPU training performance as a basis to calculate the scalability speedups.
-4.2.1. Single-host scalability (NVLink)
-We now report the single-host scalability of our two approaches by varying the number of GPUs. We consider the GAT model since it is the most computationally demanding among the ones we considered.
-Table 8 presents the results from our scalability experiment.
-The performance trends when training on a single GPU are similar to the ones we observed for the multi-GPU setups we discussed previously.
-The full-graph training systems have a much lower epoch time than the mini-batch ones but a higher time-to-accuracy.
+DistDGL 1 9.83 1.0x 11.56 80.93 1.0x
+2 7.31 1.3x 10.75 75.24 1.1x
+3 6.82 2.0x 6.61 46.28 1.7x
+Quiver 1 7.33 1.0x 2.63 18.41 1.0x
+2 4.67 1.6x 1.96 13.72 1.3x
+3 3.78 1.9x 1.66 11.62 1.6x
+Next, we answer Q2: How do systems scale with a varying number of GPUs and a varying number of hosts?
+For each dataset, we pick a GNN architecture that fits in a single GPU and use the single GPU training performance to calculate the scalability speedups.
+Single-host scalability (NVLink).
+We now measure single-host scalability by varying the number of GPUs.
+We report the hyperparameters in the appendix in table 14 For the orkut dataset, we do not report TTA since it has synthetic input features and target labels.
+Table 5 presents the results for GAT models. Like we observed previously, the full-graph training systems have a much lower ET than the mini-batch ones but a higher TTA in all configurations.
+Mini-batch systems have a lower TTA than full-graph ones also in a single GPU setting, indicating that they have a lower computation cost to start with.
+The mini-batch systems scale almost linearly, up to 3.9×3.9$\times$3.9 ×.
 When scaling to multiple GPUs, one important advantage of mini-batch training systems is that they can keep the amount of work per GPU constant by increasing the mini-batch size.
 This is not possible in full-graph training systems since the total amount of work per epoch is constant.
-As the number of GPU increases, the amount of work per GPU decreases but the communication cost increases, making it difficult to achieve scalability.
-This effect is clear when comparing systems using synchronous training.
-With PipeGCN, which uses full-graph training, the cost of synchronous cross-GPU communication during the forward and backward pass of training hides the gains of reducing the computation cost per GPU.
-Several optimizations in full-graph training systems address this problem.
-PipeGCN has a much better scalability when using asynchronous training.
-This enables PipeGCN to use pipelining, one of its most important features, to overlap communication and computation and to scale to up to 3.4×3.4$\times$3.4 ×.
+As the number of GPUs increases, the amount of work per GPU decreases but the communication cost increases, hindering scalability.
+This explains why full-graph training systems that optimize communication scale better than the Full-Graph baseline.
+PipeGCN has much better scalability because it overlaps communication and computation using asynchronous training, scaling to up to 3.4×3.4$\times$3.4 ×.
 Sampling of boundary nodes in BNS-GCN also enables good scaling, up to 3.5×3.5$\times$3.5 ×, since the cross-GPU communication is avoided.
-The mini-batch systems scale almost linearly despite using synchronous training: both DGL and Quiver scale by up to 3.9×3.9$\times$3.9 ×.
-Besides scalability, these systems achieve lower time-to-accuracy that the full-graph training systems in all settings.
-We only show ET and ET Speed-up for orkut dataset, since it is a SNAP dataset (Yang and Leskovec, 2012) with only graph structure and synthetic input features and target labels.
-We also experiment with GraphSAGE model, and the results are shown in the extended version (Bajaj et al., 2024).
-For GraphSAGE we see that synchronous PipeGCN doesn’t scale at all, asynchronous PipeGCN scales up to $\times3$ and BNS-GCN scales up to 2.7×2.7$\times$2.7 ×.
-While DGL and Quiver are able to scale up to 4.0$\approx$ 4.0$\times$ 4.0 × and 3.8×3.8$\times$3.8 × respectively.
+Results on GraphSAGE models are shown in the appendix in table 21 The Full-Graph baseline does not scale at all, PipeGCN scales up to $\times3$ and BNS-GCN scales up to 2.7×2.7$\times$2.7 ×.
+In contrast, DGL and Quiver are able to scale up to 4.0$\approx$ 4.0$\times$ 4.0 × and 3.8×3.8$\times$3.8 × respectively.
+Multi-host scalability (NVLink).
+We now consider scaling the training of larger graphs (orkut and ogbn-papers100m) in a distributed setting of NVLink hosts, where each host has 4 GPUs.
+For this experiment, we replace DGL with DistDGL, its distributed version.
+We report the hyperparameters in the appendix in table 15.
+We consider the GraphSAGE model because it is the only one implemented in the DistDGL distribution.
+We run Quiver with data parallelism across multiple hosts.
+Figure 4. Convergence curve for GraphSage (3 NVLink hosts).
+Table 6 and Figure 4 show the results for distributed scalability with NVLink-enabled devices.
+Also, in this case, we observe that full-graph training systems consistently have lower epoch time but larger time-to-accuracy compared to DistDGL.
+Going from single-host to multi-host training impacts different systems in a different way.
+Full-graph training systems must use the network to exchange vertex features across hosts, which is much slower than NVLink.
+The Full-Graph baseline actually has a slower epoch time than a single-host setup.
+Optimizing communication using asynchrony (PipeGCN) and boundary node sampling (BNS-GCN) is even more crucial to scaling than in the single-host scalability experiments.
+DistDGL partitions the dataset across multiple hosts and performs distributed sampling, which introduces an additional communication cost compared to a single-host implementation.
+When running on multiple hosts, the DistDGL baseline has a much larger epoch time than Quiver because it uses distributed CPU-based sampling and does not use caching.
+Scalability with the PCIe host.
+We run single-host scalability experiments on the PCIe host, from 1 to 4 GPUs.
+Like for the previous experiments, mini-batch training systems have a lower time-to-accuracy than full-graph training ones despite having a larger epoch time.
+The results show that all the full-graph training systems have lower scalability since they must not communicate over the slower PCIe bus instead of NVLink.
+This slowdown compared to the NVLink case is particularly evident for the Full-Graph baseline, which can only scale up to 2.0×2.0$\times$2.0 ×.
+PipeGCN async can scale up to at most 3.1×3.1$\times$3.1 × and BNS-GCN up to at most 3.0×3.0$\times$3.0 ×.
+Mini-batch training systems have better scalability because they replicate most graphs in the cache and require less cross-GPU communication.
+DGL scales up to 3.8×3.8$\times$3.8 × and Quiver up to 3.5×3.5$\times$3.5 ×.
+We also run all systems using 3 PCIe hosts for the ogbn-papers100m graph.
+Compared to the distributed NVLink setup, hosts are connected by a much slower network (1 Gbps instead of 25 Gbps).
+Like in the previous experiments, DistDGL has a lower time-to-accuracy than the other full-graph training systems, but the difference is smaller than with the distributed NVLink setup because of the cost of distributed sampling and feature loading over a very slow network.
+DistDGL outperforms the BNS-DGL system by only $12\%$.
+The distributed sampling implementation of DistDGL is not optimized for very slow network links.
+The full results are shown in the appendix in table 20.
+Memory scalability.
+In terms of memory usage, full-graph training systems must keep the entire graph and the related training state partitioned in GPU memory.
+Mini-batch systems only need to store a micro-batch and the related state in the memory of each GPU.
+In our experiments, full-graph training systems have a larger peak memory utilization in all the configurations we considered, up to 7.4×7.4$\times$7.4 ×, but the gap decreases when using more GPUs (since mini-batch training systems must store more micro-batches in total), deeper GNNs with more layers, or larger sampling fanouts (since each micro-batch is larger).
 Takeaway.
-The mini-batch training systems show better scalability than the full-graph training ones.
-This is because they can tune the mini-batch size to keep the work per GPU constant to maintain high GPU utilization, unlike the full-graph training systems.
-The full-graph training systems also experience increasing communication costs as the number of GPUs increases.
-However, these overheads can be mitigated using asynchronous training and boundary nodes sampling.
-4.2.2. Distributed scalability (NVLink)
-We now consider scaling the training of larger graphs (orkut and ogbn-papers100m) in a distributed setting of NVLink hosts, where each host has four GPUs.
-We consider the GraphSAGE model because it is the only one implemented in the DistDGL distribution.
-We do not consider Quiver because it does not support distributed training.
-Table 9 shows the results for distributed scalability with NVLink enabled devices.
-The synchronous version of PipeGCN cannot scale when increasing the number of hosts.
-Using asynchronous training and boundary nodes sampling can make full-graph training systems scale much better, as we also observed for the single-host experiments.
-DistDGL has similar scalability trends despite the fact that it does not use GPU caching.
-Despite its higher epoch time compared to full-graph training systems, DistDGL achieves a much lower time-to-accuracy.
-Takeaway.
-The general scalability considerations we made for single-host scalability apply also to multi-host scalability.
-Synchronous full-graph training does not scale to multiple host.
-By using asynchronous training and boundary nodes sampling, PipeGCN and BNS-GCN achieve similar scalability as DistDGL, which uses mini-batch training.
-However, DistDGL has a lower time-to-accuracy in all configurations.
-4.2.3. Single-host scalability (PCIe hosts)
-We run the same experiments on PCIe hosts (as done in Section 4.1.3), and observe similar results, which are shown in the. We scale the number of GPUs from one to four.
-All the full-graph training systems have a lower scalability because they require communication over the PCIe bus, which is slower than NVLink.
-This slowdown compared to the NVLink case is particularly evident for PipeGCN sync, which can only scale up to 2.0×2.0$\times$2.0 ×.
-PipeGCN async can scale up to at most 3.1×3.1$\times$3.1 × and BNS-GCN up to at most 3.0×3.0$\times$3.0 ×.
-Mini-batch training systems have a better scalability because they require little cross-GPU communication. DGL scales up to 3.8×3.8$\times$3.8 × and Quiver up to 3.5×3.5$\times$3.5 ×.
-Takeaway.
-The full-graph systems are more sensitive to the performance of the cross-GPU interconnect because they need to exchange vertex features between partitions at each GNN layer during the forward and backward pass.
-When GPUs must communicate over a slow PCIe interconnect, the scalability of these systems is lower.
-The mini-batch training systems use data-parallelism, which does not require exchanging vertex features across GPUs.
-Therefore, these systems show better scalability in the PCIe host.
+Mini-batch training systems have lower time-to-accuracy than full-graph training ones for all models, datasets, and hardware configurations we consider.
+They also have better scalability in most cases.
+One important scalability advantage for mini-batch training systems is that they can tune the mini-batch size to keep the work per GPU constant.
+In full-graph training systems, algorithmic optimizations such as asynchrony and sampling are essential to achieve high scalability.
+The full-graph systems cannot scale well when they use slower cross GPU interconnects because their communication cost is amplified.
 5. Accuracy Evaluation
-Even though a system takes longer than another to reach the same target accuracy, it could still be preferable to use it if it can converge to a higher accuracy.
-Indeed, some works on GNN systems assert that full-graph training achieves higher accuracy and that mini-batch training does not guarantee convergence (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
-Figure 4 summarizes the accuracy values reported in the literature we reviewed for different GNN models and datasets.
+Even though a system takes longer than another to reach the same target test accuracy, it could still be preferable to use it if it can converge to a higher test accuracy.
+Indeed, some works on GNN systems assert that full-graph training achieves higher accuracy and that mini-batch training does not guarantee convergence (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
+Figure 5 summarizes the test accuracy values reported in the literature we reviewed for different GNN models and datasets.
 Each point in the graph represents the test accuracy reported for a specific GNN model-dataset pair in one of the papers we reviewed.
 The figure illustrates a wide distribution of accuracy values for the same dataset and model combination.
-In this Section, we shed light onto this apparent contradiction and answer two questions: Q3) What is the best accuracy we can achieve using full-graph and mini-batch training?
-Q4) What accuracy can we achieve when using a training method with the best hyperparameters found for the other method?
-In practice, the cost of training a model also depends on how long it takes to find good hyperparameters.
-Therefore, we also consider the following question: Q5) How sensitive are the two training methods to changes in the hyperparameters?
+In this Section, we shed light onto this apparent contradiction and answer two questions: Q3) What is the test best accuracy we can achieve using full-graph and mini-batch training?
+Q4) What test accuracy can we achieve when using a training method with the best hyperparameters found for the other method?
+Different training systems have different optimizations, so Q5: What is the impact of the optimizations on accuracy?
 Figure 5.
-Test accuracy reported in literature on GNN training.
-(Wan et al., 2022b; dgl, 2023; Hu et al., 2020; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Md et al., 2021; Yang et al., 2023; Li et al., 2024; Gasteiger et al., 2022; Mostafa, 2022; Kaler et al., 2022; Yao and Li, 2021; Zou et al., 2019; Hamilton et al., 2017; Chiang et al., 2019; Zeng et al., 2019; Chen et al., 2018b, a; Yao and Li, 2021; Wu et al., 2024; Veličković et al., 2018; Thorpe et al., 2021; Frasca et al., 2020; Wan et al., 2022a; Kipf and Welling, 2017; Jia et al., 2020)
-Methodology To compare maximum accuracies among full-graph and mini-batch training, we use the methodology depicted in Figure 5.
+Test accuracy reported in literature on GNN training (Wan et al., 2022b; dgl, 2023; Hu et al., 2020; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Md et al., 2021; Yang et al., 2023; Li et al., 2024; Gasteiger et al., 2022; Yao and Li, 2021; Zou et al., 2019; Hamilton et al., 2017; Chiang et al., 2019; Zeng et al., 2019; Chen et al., 2018b; Yao and Li, 2021; Wu et al., 2024; Veličković et al., 2018; Thorpe et al., 2021; Frasca et al., 2020; Wan et al., 2022a; Kipf and Welling, 2017; Jia et al., 2020) .
+Methodology.
+To compare maximum test accuracies among full-graph and mini-batch training, we use the methodology depicted in Figure 6.
 Given a dataset and a GNN model, we conduct two separate hyperparameter searches, one using full-graph training and one using mini-batch training.
-This is because the two training methods can have different sets of best hyperparameters for the same dataset and model.
-The space of hyperparameters we consider is shown in Table 3.
+This is because the two training methods can have different sets of best hyperparameters for the same dataset and model, using the method described in Section 3.
+We consider vanilla baselines for full-graph and mini-batch training: the Full-Graph baseline and DGL with the default Neighborhood Sampling algorithm, respectively.
 The best hyperparameter settings we obtained using the two training methods are denoted as FG-search and MB-search.
 We then run full-graph and mini-batch training on both settings, measure the test accuracy, and denote it as FG-train and MB-train respectively.
-This yields a total of four accuracy combinations per dataset and GNN model.
-We report all the hyperparameter settings we found in the extended version of this paper (Bajaj et al., 2024).
-To factor out any accuracy bias due to systems-level optimizations, we perform synchronous training with PipeGCN and DGL.
+This yields a total of four test accuracy combinations per dataset and GNN model.
+We report the hyperparameter settings we found in the appendix in table 12.
 We validate our results by verifying that given a dataset, model, and training method, our hyperparameter search achieves accuracy values that are equal to or better than the best accuracy value reported in the literature reviewed in Section 2.3.
 Figure 6. Methodology of our accuracy comparison between full-graph (FG) and mini-batch (MB) training.
-5.1. Accuracy with the Best Setting
-We start by answering question Q3) What is the best accuracy we can achieve using full-graph and mini-batch training?
-Answering this question shows whether there is one method that consistently reach higher accuracy than another.
-When the hyperparameters are properly set, whether full-graph training achieves higher accuracy than mini-batch training or vice versa depends on the dataset-GNN model pairs.
+Table 7. Accuracy (%) of GraphSage using hyperparameter tuning method in Figure 6.
+Pubmed Ogbn-arxiv Reddit Ogbn-products Ogbn-papers100M
+FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search
+FG-train $\textbf{77.40}\pm\textbf{00.04}$ $76.10\pm 00.67$ $\textbf{71.76}\pm\textbf{00.16}$ $67.21\pm 00.9$ $\textbf{96.94}\pm\textbf{00.01}$ $90.28\pm 00.63$ $\textbf{79.29}\pm\textbf{00.15}$ $76.60\pm 00.32$ $\textbf{62.82}\pm\textbf{00.01}$ $61.74\pm 00.13$
+MB-train $77.00\pm 00.39$ $\textbf{78.20}\pm\textbf{00.32}$ $70.73\pm 00.38$ $\textbf{71.67}\pm\textbf{00.30}$ $96.52\pm 00.14$ $\textbf{96.82}\pm\textbf{00.04}$ $78.72\pm 00.07$ $\textbf{78.84}\pm\textbf{00.12}$ $62.52\pm 00.04$ $\textbf{62.77}\pm\textbf{00.02}$
+Table 8. Accuracy (%) impact of different optimizations - GraphSAGE.
+Pubmed Ogbn-arxiv Reddit Ogbn-products Ogbn-papers100M
+Full-Graph $77.40\pm 0.04$ $71.76\pm 0.16$ $96.94\pm 0.01$ $79.29\pm 0.15$ $62.82\pm 0.01$
+PipeGCN $77.40\pm 0.45$ $71.27\pm 0.66$ $96.93\pm 0.11$ $78.58\pm 0.32$ $61.74\pm 0.13$
+BNS-GCN $77.20\pm 0.16$ $71.72\pm 0.27$ $96.95\pm 0.08$ $78.40\pm 0.22$ $62.79\pm 0.10$
+AdaQP $77.42\pm 0.59$ $63.73\pm 0.07$ $96.85\pm 0.002$ $78.79\pm 0.004$ $62.54\pm 0.00$
+NS $78.20\pm 0.32$ $71.67\pm 0.30$ $96.82\pm 0.04$ $78.84\pm 0.12$ $62.77\pm 0.02$
+ClusterGCN $79.72\pm 0.22$ $65.72\pm 0.18$ $95.07\pm 0.03$ $79.50\pm 0.17$ $49.79\pm 0.00$
+GraphSaint $84.03\pm 0.61$ $73.80\pm 0.13$ $98.40\pm 0.03$ $77.85\pm 0.07$ $50.32\pm 0.27$
+Accuracy with the best setting.
+We start by answering question Q3: What is the best test accuracy we can achieve using vanilla full-graph and mini-batch training?
 The accuracy results for the GraphSAGE model are shown in Table 7.
@@
 For the GraphSAGE model, each method can yield higher accuracy than the other, depending on the dataset.
+Accuracy tends to have lower variance with full-graph training generally than with mini-batch training.
 The largest gap between the accuracies obtained with the two methods is less than $0.9\%$ in favor of full-graph training and averaging around $0.3\%$.
-We ran a similar analysis for the GCN and GAT model, and we report the results in the extended version of this paper (Bajaj et al., 2024).
+We ran a similar analysis for the GCN and GAT model, and we report the results in the appendix in table 18 and table 19.
 With the GAT model, the largest gap is slightly higher, up to $1.1\%$ in favor of full-graph training and the average is around $0.6\%$.
@@
 The largest gap in this case is $3.6\%$ in favor of mini-batch training and the average is around $1.27\%$.
-Takeaway.
-These results indicate that no training method alone guarantees higher accuracy than the other and the gap between the two methods is typically small (i.e., less than 3.6% accuracy).
-If achieving the highest accuracy is an absolute priority regardless of the training cost, it is advisable to conduct hyperparameter search twice using both training methods.
-Otherwise, using either of the two methods is likely to produce similar accuracy in the absence of other systems-level optimizations such as asynchronous training or boundary node sampling.
-Table 10. Accuracy (%) of GraphSage using hyperparameter tuning method in Figure 5.
-Pubmed Ogbn-arxiv Reddit Ogbn-products Ogbn-papers100M
-FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search
-FG-train $\textbf{77.40}\pm\textbf{00.04}$ $76.10\pm 00.67$ $\textbf{71.76}\pm\textbf{00.16}$ $67.21\pm 00.9$ $\textbf{96.94}\pm\textbf{00.01}$ $90.28\pm 00.63$ $\textbf{79.29}\pm\textbf{00.15}$ $76.60\pm 00.32$ $\textbf{62.82}\pm\textbf{00.01}$ $61.74\pm 00.13$
-MB-train $77.00\pm 00.39$ $\textbf{78.20}\pm\textbf{00.32}$ $70.73\pm 00.38$ $\textbf{71.67}\pm\textbf{00.30}$ $96.52\pm 00.14$ $\textbf{96.82}\pm\textbf{00.04}$ $78.72\pm 00.07$ $\textbf{78.84}\pm\textbf{00.12}$ $62.52\pm 00.04$ $\textbf{62.77}\pm\textbf{00.02}$
-5.2. Accuracy with Sub-Optimal Setting
-We now answer the next question Q4) What accuracy can we achieve when using a training method with the best hyperparameters found for the other method?
-This shows how the outcome of the comparison between the two training approaches can change when using sub-optimal hyperparameter settings.
+Accuracy with sub-optimal setting. We now answer question Q4: What test accuracy can we achieve when using a vanilla training method with the best hyperparameters found for the other method?
 To answer this question for the case of the GraphSAGE model, we still look at Table 7.
@@
 Full-graph training has lower accuracy than mini-batch training and the gap between the two methods is larger, reaching $6.22\%$ for the reddit dataset.
-Our evaluation on GAT and GCN, which is reported on the extended version (Bajaj et al., 2024), leads to similar conclusions.
-For GAT, largest difference between FG-train and MB-train is $2.85\%$ in favor of MB-search for the ogbn-arxiv dataset and $1.9\%$ in favor of FG-search for the pubmed dataset.
-For GCN, it is $1.36\%$ in favor of FG-search for the reddit dataset and $5.5\%$ in favor of MB-search for the pubmed dataset.
+Our evaluation of GAT and GCN is reported on the appendix in table 9 and table 10 which leads to similar conclusions.
+For GAT, the largest difference between FG-train and MB-train is $2.85\%$ in favor of MB-search for arxiv and $1.9\%$ in favor of FG-search for pubmed.
+For GCN, it is $1.36\%$ in favor of FG-search for reddit and $5.5\%$ in favor of MB-search for pubmed.
+These results motivate the use of separate hyperparameter tuning for the two methods since good hyperparameter settings do not transfer well across methods.
+Impact of optimizations on accuracy.
+Many performance optimizations we consider can result in a different accuracy than the vanilla methods.
+This section answers the question Q5: What is the impact of the optimizations on accuracy?
+In this analysis, we use the best-found hyperparameters for each system class: FG-Search and MB-Search models for full-graph and mini-batch training systems, respectively.
+For full-graph training, we consider the impact of asynchrony (PipeGCN), sampling (BNS-GCN), and message quantization (AdaQP).
+For mini-batch training, besides Neighborhood Sampling (NS), we also consider the ClusterGCN and GraphSaint sampling algorithms.
+The results for GraphSage are shown in Table 8.
+Among full-graph training systems, our results show that asynchrony and sampling have minimal impact on accuracy compared to FG, within 1% in all the cases we considered.
+Message quantization has a stronger negative impact only on arxiv.
+Mini-batch training shows a very different trend.
+For most datasets, there is some sampling algorithm that achieves higher accuracy than all full-graph training approaches.
+The relative performance of each sampling algorithm shows large variations depending on the dataset.
+Both GraphSaint and ClusterGCN achieve the best and the worst accuracy among all sampling algorithms for some datasets.
+These results show that the regularization effect of sampling is key to achieving high accuracy, focusing the training on nodes that have the highest influence on each other (Zeng et al., 2019).
+Neighborhood Sampling (NS) shows a more stable behavior: on all datasets, it achieves similar accuracy as FG.
+This justifies its popularity as the default sampling method for mini-batch training.
+We repeated the experiment on GAT and GCN and obtained similar trends, as shown in the appendix in table 18 and table 19.
 Takeaway.
-These results show that given a dataset and GNN model, each training method can outperform the other depending on the choice of the hyperparameters.
-This corroborates our choice of performing two separate hyperparameter searches, one per training method, for our comparison.
-It also potentially explains why existing work shows inconsistent evidence about which training method leads to higher accuracy.
-We hypothesize that the large variations in accuracy shown in Figure 4 is due to the difference in the choice of hyperparameter settings, which can greatly impact the accuracy and performance of GNN training.
-Mini-batch training still performs relatively well in settings where full-graph training performs best, but the opposite is not true.
-This is an indication that mini-batch training is more robust to different hyperparameter settings, potentially making hyperparameter search easier.
-We will further corroborate this thesis in later experiments.
-5.3. Ease of Hyperparameter Search
-Figure 6. Sensitivity to the number of layers and hidden size for GraphSAGE model.
-We finally answer our last question focusing on accuracy and hyperparameter search, which is Q5) How sensitive are the two training methods to changes in the hyperparameters?
-We evaluate the ease of finding a good hyperparameter setting using the two training approaches by measuring how accuracy varies as we vary the model architecture hyperparameters, i.e., the number of hidden layers and the size of the hidden vertex dimensions.
-We trained each model for a maximum of 1000 epochs, stopping earlier if we the validation accuracy doesn’t increase for 50 epochs.
-Figure 6 illustrates the best test accuracy when training the GraphSAGE model, top row shows mini-batch training experiments and the bottom row shows synchronous full-graph training.
-Notably, mini-batch training demonstrates remarkable consistency across varying model architectures.
-In contrast, full-graph training exhibits a wider range of variations in test accuracy when the architecture changes.
-We observed similar results for GCN and GAT models, which are shown in the extended version of this paper (Bajaj et al., 2024).
-Figure 6 and the other results for GCN and GAT, which are reported in the extended version (Bajaj et al., 2024), clearly show the effect of over-smoothing, which appears in GNNs at increased depths (Rusch et al., 2023).
-Mini-batch training on GraphSAGE does not show smoothing with up to 8 layers in Figure 6.
-In a set of separate experiments we found that oversmoothing does occur with mini-batch training, but only after we increase the number of layers to $16$ or higher.
-Our experiments for GCN and GAT show that oversmoothing occurs at lower depths for mini-batch training but the resulting accuracy drop is still much less than that with full-graph training (see the extended version (Bajaj et al., 2024)).
+When comparing vanilla full-graph and mini-batch training, no method consistently guarantees higher accuracy than the other and the gap between the two methods is not larger than 3.6%.
+Good hyperparameter settings do not perform as well across both full-graph and mini-batch training, motivating the need for separate tuning processes.
+The algorithmic optimizations for full-graph training we considered achieve very similar accuracies as the vanilla full-graph method in most cases, making them attractive choices given their performance benefits.
+For mini-batch training, using the right sampling algorithm can achieve higher accuracy than the full-graph training methods for most datasets and models.
+However, some sampling algorithms show widely different accuracies based on the dataset and sometimes converge to much lower accuracies.
+This indicates that trying different algorithms during hyperparameter tuning is essential.
+Overall, these results challenge the common rationale for using full-graph training over mini-batch training, suggesting that the expected benefits of avoiding sampling and its associated information loss to achieve higher accuracy do not materialize in practice.
+6. Cost Analysis
+To support and generalize our empirical observations beyond the specific hardware configurations and software implementations we evaluated empirically, in this section we answer the following question Q6: What are the analytical performance costs of vanilla full-graph and mini-batch training?
+We model the communication and computation costs of training analytically and then evaluate the cost of pre-processing (sampling) in mini-batch training experimentally to complement the analysis.
+Communication cost.
+Our analysis considers a simplified model of GNN training where the training workload is partitioned among a set of workers $W$, each accessing a local memory.
+The communication cost is the volume of vertex feature data exchanged among the workers.
+We ignore the communication cost of gradient synchronization since GNN models have relatively few parameters.
+We measure the communication cost to convergence rather than the cost per epoch to account following the discussion of Section 4.1.
+In full-graph training, at each layer, workers must receive the features of all the vertices in other partitions that have neighbors in the local partition. The communication cost is:
+$\Gamma_{\textrm{fg}}=n_{f}\sum_{l=1}^{L}\sum_{w\in W}\sum_{v\in R_{w}}|h_{v}^{l}|$
+where $n_{f}$ is the number of epochs to convergence, $R_{w}$ the set of vertices that are remote neighbors of vertices in $w$, $h_{v}^{l}$ is the feature vector for vertex $v$ at layer $l$.
+In mini-batch training, we model a scenario where the input features cannot fit in the local memory of one worker and are partitioned among workers.
+At each iteration, each worker must gather the input features of all the vertices at the bottom layer of its micro-batch.
+The communication cost to convergence is:
+$\Gamma_{\textrm{mb}}=n_{m}\sum_{i=1}^{I}\sum_{w\in W}\sum_{v\in(M_{i,w}\setminus P_{w})}|h_{v}^{0}|$
+where $n_{m}$ is the number of epochs to convergence using vanilla mini-batch training, $I$ is the number of iterations in an epoch, $M_{i,w}$ is the set of vertices in the micro-batch assigned to worker $w$ at iteration $i$, and $P_{w}$ is the partition of vertices assigned to worker $w$.
+We calculate the ratio between $\Gamma_{\textrm{fg}}$ and $\Gamma_{\textrm{mb}}$ by considering the same hyperparameters for both approaches, which are the same we used for evaluating the time-to-accuracy in Section 4.1.
+We use Metis to partition the datasets (Karypis and Kumar, 1997) and run Neighborhood Sampling for one epoch to obtain $M^{i}_{w}$.
+Figure 7 shows the ratio between $\Gamma_{\textrm{fg}}$ and $\Gamma_{\textrm{mb}}$ for different datasets.
+The communication cost of full-graph training is higher than that of mini-batch training in most cases.
+The gap mostly depends on the number of boundary nodes and, for mini-batch training, on the size of the last layer of the micro-batches.
+In practice, mini-batch training does not need to partition the dataset since it can replicate commonly accessed input features in the local memory of multiple workers.
+In our experimental evaluation of Section 4, all datasets except orkut and papers100M are fully replicated, but partial replication is often almost as effective (Song et al., 2023b).
+In general, larger graphs can have a larger edge cut when they are partitioned.
+This results in larger $R_{w}$, increasing the communication cost of full-graph training $\Gamma_{\textrm{fg}}$.
+In contrast, the communication cost of mini-batch training, $\Gamma_{\textrm{mb}}$, grows with $(M_{i,w}\setminus P_{w})$, which is upper bounded by the size of the sampled micro-batches, not by the size of the graph.
+A similar effect is observed by increasing the number of partitions with fixed datasets, as shown in Figure 7.
+This increases the relative cost of $\Gamma_{\textrm{fg}}$ over $\Gamma_{\textrm{mb}}$.
+Figure 7. FG/MB ratio of communication cost. (Left: Hyperparameters in the appendix in table 16. Right: Default 4 partitions)
+Figure 8. FG/MB ratio of computation cost for GraphSage.
+Computation cost. We now analyze the computational cost of training with vanilla full-graph and mini-batch training. We consider the same model used for the communication cost analysis.
+A GNN training layer $l$ computes the features of vertices at layer $l$ based on the features of their neighbors at layer $l-1$ (see Section 2.1).
+The computational cost to the convergence of full-graph training can be expressed as:
+$\Theta_{\textrm{fg}}=n_{f}\sum_{l=1}^{L}(|E^{l}|\cdot c_{e}+|V^{l}|\cdot c_{v})(1+\eta)$
+where the summation expresses the cost of the forward pass: $c_{e}$ is the cost of generating and aggregating a message sent over one edge, $c_{v}$ is the cost of computing a new representation of a vertex, $V^{l}$ is the set of all vertices processed at layer $l$ in the epoch, and $E^{l}$ is the set of all edges between $V^{l-1}$ and $V^{l}$ processed in the epoch.
+The factor $(1+\eta)$ accounts for the cost of the backward pass.
+The values of $c_{e}$ and $c_{v}$ are independent of the choice of full-graph or mini-batch training.
+They depend on the GNN model (e.g.
+GraphSage or GAT) and its hyperparameters and they can be calculated analytically in terms of FLOPS.
+The computational cost for mini-batch training is similar:
+$\Theta_{\textrm{mb}}=n_{m}\sum_{i=1}^{I}\sum_{w\in W}\sum_{l=1}^{L}(|E^{l}(M_{i,w})|\cdot c_{e}+|V^{l}(M_{i,w})|\cdot c_{v})(1+\eta)$
+where $E^{l}(M_{i,w})$ and $V^{l}(M_{i,w})$ are the number of edges and vertices at layer $l$ in the micro-batch $M_{i,w}$.
+We calculate the ratio of $\Theta_{\textrm{fg}}$ and $\Theta_{\textrm{mb}}$ using the same hyperparameters as in the communication cost analysis and show it in Figure 8.
+The results reveal a large difference in terms of computational cost between mini-batch and full-graph training, especially with the largest graph, papers100M.
+In full-graph training, $\Theta_{\textrm{fg}}$ depends on $|E^{l}|$ and $|V^{l}|$, which grow as the size of the graph grows.
+Mini-batch training performs message passing only on a sample of the graph, unlike full-graph training, reducing computation cost substantially.
+The computation cost $\Theta_{\textrm{mb}}$ is upper-bounded by the size of $M_{i,w}$, which depends on the number of layers in the GNN and the sampling algorithm rather than on the size of the graph.
+As the number of workers grows, the gap between full-graph and mini-batch training decreases in Figure 8.
+This is because the same vertex can appear in multiple micro-batches, so its hidden features are computed multiple times in an epoch.
+This redundant computation does not happen with full-graph training.
+Cost of sampling in mini-batch training.
+Mini-batch training requires two pre-processing steps: sampling and data loading.
+Data loading costs are included in the communication costs.
+Modeling sampling costs in terms of FLOPS, as we did for the training, costs $\Theta_{\textrm{mb}}$ is challenging because sampling is a sparse and irregular computation, not a dense computation like matrix multiplication.
+Therefore, to put the previous results in context, it is necessary to consider the relative computational cost of sampling over training in mini-batch training.
+We measure it experimentally using DGL and Neighborhood Sampling, running both sampling and training in CPU memory to factor out the overhead of transferring data to the GPUs.
+The results are shown in Figure 9.
+The computational cost of sampling is noticeable but is still lower than the training cost.
+It decreases when we increase the number of workers/partitions because each epoch requires fewer iterations.
 Takeaway.
-These results highlight the robustness of mini-batch training to alterations in model architecture.
-Full-graph training appears to be more sensitive to changes in architecture, so it requires a more careful and potentially expensive hyperparameter search.
-Whereas mini-batch training is consistently less sensitive to changes of those parameters than full-graph training, across all the GNN models we consider.
-In addition, mini-batch training is much less prone to the oversmoothing problem (Rusch et al., 2023) that affects deep GNN models with many layers.
+This analysis shows that vanilla full-graph training has a higher communication cost than vanilla mini-batch training, especially with larger graphs.
+In practice, mini-batch training can replicate input features across multiple workers to eliminate or reduce communication costs.
+The high communication cost of full-graph training has been the focus of prior work on full-graph training systems, as discussed in Section 2.2.
+The analysis shows that using sampling in mini-batch training also results in a substantial reduction in computation cost compared to full-graph training systems, especially with larger graphs.
+Overcoming this gap while maintaining the full-graph training approach and avoiding sampling is challenging.
+Figure 9. Fraction of epoch time spent performing sampling in mini-batch training for GraphSage.
 7. Conclusion
-This work is motivated by the need to understand the state of the art in GNN training systems, given the absence of consistent empirical evidence comparing full-graph and mini-batch training systems.
-It first conducted an in-depth literature survey and investigated the reasons for the inconsistent claims and the lack of empirical evidence in the literature.
-It further provided a thorough empirical evaluation comparing representative systems supporting the two approaches on various GNNs and datasets.
-The study reveals that the mini-batch training systems we considered consistently achieve faster convergence time, measured as time-to-accuracy, compared to full-graph training, regardless of the choice of the dataset, GNN models, and hardware settings.
-The study also shows that using time-to-accuracy as a metric is more relevant than comparing epoch times for systems that have different convergence criteria.
-Systems of both types can achieve similar accuracy when hyperparameter tuning is performed separately for full-graph and mini-batch training.
-Our work stresses the importance of comparing GNN training systems of both classes in future work in this area.
-Using time-to-accuracy as the main metric for comparison across areas is important, as is performing separate hyperparameter tuning for the two training methods.
-These findings are relevant to both machine learning practitioners who utilize GNN training systems and systems researchers involved in their design and evaluation.
+A common rationale for using full-graph training is to achieve high accuracy by avoiding the information loss results from sampling.
+In our evaluation, we did not observe these expected benefits.
+Mini-batch training systems not only achieve a lower time-to-accuracy across all models, datasets, and hardware configurations we consider but also a comparable or even higher accuracy with proper hyperparameter tuning.
+Mini-batch training also has some limitations.
+The common Neighborhood Sampling algorithm is able to consistently achieve good accuracy, but it still shows a slightly larger variance between training runs than vanilla full-graph training.
+Other sampling algorithms we evaluated are not always able to converge to a high accuracy, depending on the dataset.
+Finding sampling algorithms that have consistently higher performance than Neighborhood Sampling is an interesting avenue of research.
+Our experiments show that as GNN models become deeper, the advantage of mini-batch training over full-graph training reduces.
+Distributed sampling over multiple hosts can also become a bottleneck, especially when the network is slow.
+An important advantage of mini-batch training systems is that the sampling algorithms are interchangeable components.
+This modularity and flexibility in choosing the right sampling algorithm is key since different algorithms perform very differently on different datasets.
+Our experiments show that providing efficient sampling implementations is important for multi-host setups, especially when using a slow network.
+In order to support future innovations in sampling algorithms, it is important to have systems that allow ML practitioners to implement new sampling algorithms and run them on GPUs, such as (Jangda et al., 2021; Pandey et al., 2020; Gong et al., 2023; Wang et al., 2021).
+In terms of improving full-graph training systems, our cost model analysis and single-GPU performance experiments reveal that in addition to the well-known communication cost of full-graph training—a bottleneck addressed by much previous work—there is a computation cost gap that is difficult to bridge without employing sampling.
+A natural open question is whether combining full-batch training with more aggressive sampling, forgoing full-graph aggregation, can be more effective than standard mini-batch training.
+In our hyperparameter tuning, however, we found that the accuracy of mini-batching typically peaks with batch sizes around 1024/4096 and does not grow further with larger values, so the advantages of using full-batch training are unclear.
 Appendix
-The results for highest achievable accuracy for GAT and GCN model are shown in Table 11 and Table 12.
+The results for the highest achievable accuracy for the GAT and GCN model are shown in Table 9 and Table 10.
 Given a dataset, we must compare the accuracies obtained for two combinations: (FG-search, FG-train) for full-graph training and (MB-search, MB-train) for mini-batch training.
 The results are averaged after 5 runs.
+Table 11 shows the search space of our hyperparameter search.
 Table 12 shows the best architectures obtained for every model-dataset combination.
+Table 13 shows the architectures used for TTA comparisons.
+Tables 14 and  15 show the architectures used for scalability comparisons.
+Table 16 shows the architectures used for communication cost experiments.
 Figure 11 and Figure 12 illustrate the best test accuracy when training the GCN and GAT respectively.
@@
 MB-train $78.20\pm 00.16$ $78.70\pm 00.23$ $70.54\pm 00.16$ $71.00\pm 00.16$ $93.31\pm 00.14$ $94.41\pm 00.14$ $75.27\pm 00.07$ $79.31\pm 00.04$
+Table 11. Hyperparameter search space
+Parameter range/values
+Number of Hidden Layers [2, 10]
+Size of hidden layers [16, 1024]
+Aggregator (GraphSage only) { mean, gcn, pool }
+Number of heads (GAT only) [1, 12]
+Sampling Fanout (MB only) [4, 50]
+Learning rate [0.0001, 0.01]
+Dropout [0.2, 0.9]
+Micro-batch size (MB only) [32, 1024]
 Table 12. Best architectures
@@
 Hidden size 256 128 512 256
+Fanout 10 20 5 5
 GAT FG Layers 3 3 2 2
@@
 Num heads 4 2 2 2
+Fanout 10 10 10 5
 GCN FG Layers 2 2 2 3
@@
 Hidden size 64 1024 512 512
-Figure 7.
-Ablation study with number of layers and hidden size for GCN model.
-Legend represents the size of each hidden layer.
-X-axis shows the number of layers.
-Y-axis shows the test accuracy.
-Top row is for mini-batch training and bottom row is for full graph training.
-We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
-Figure 8.
-Ablation study with number of layers and hidden size for GAT model.
-Legend represents the size of each hidden layer.
-X-axis shows the number of layers.
-Y-axis shows the test accuracy.
-Top row is for mini-batch training and bottom row is for full graph training.
-We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
-Table 14 shows the results of the time to accuracy experiment for GAT model on a single host with NVLink connectivity.
+Fanout 10 15 15 5
+Table 13. Architectures used for TTA experiments (same for FG and MB).
+Model Architecture Pubmed Ogbn-arxiv Reddit Ogbn-products Ogbn-papers100M
+GraphSAGE Layers 3 2 4 5 2
+Hidden size 256 512 1024 256 128
+Fanout 10 20 5 5 5
+GAT Layers 3 3 2 3 -
+Hidden size 256 1024 1024 128 -
+Fanout 10 10 10 5 -
+Num Heads 4 6 4 3 -
+Table 14. Architectures used for Scalability experiments single host (same for FG and MB).
+Model Architecture Ogbn-arxiv Reddit Ogbn-products Orkut
+GAT Layers 3 2 3 3
+Hidden size 1024 1024 128 128
+Num Heads 6 4 3 2
+Fanout 25 20 20 5
+Table 15. Architectures used for Scalability experiments for 3 hosts (same for FG and MB).
+Model Architecture Orkut Ogbn-papers100m
+GraphSAGE Layers 3 2
+Hidden size 512 512
+Fanout 10 10
+Table 16. Architectures used for Communication cost experiments (Same for FG and MB).
+Model Architecture Pubmed Ogbn-arxiv Reddit Ogbn-products Ogbn-papers100M
+GraphSAGE Layers 3 2 4 5 2
+Hidden size 256 512 1024 256 128
+Fanout 10 20 5 5 5
 Table 17.
@@
 Quiver (MB, Sync) 35.96 101.68 44.61
-Table 15 shows the results for the scalability experiments for GAT model on PCIe setup and Table 16 shows the results of the scalability experiment on the NVLink setup.
+Table 17 shows the results of the time-to-accuracy experiment for the GAT model on a single host with NVLink connectivity.
+Table 18 and  19 show the accuracies achieved for GAT and GCN models for different training techniques an optimizations.
+Table 20 shows the results of the scalability experiments for the GAT model on PCIe setup and table 21 shows the results of the scalability experiment on the NVLink setup.
+Table 22 shows the results of the time-to-accuracy experiments for the GraphSAGE model on 3 PCIe hosts.
+Tables 23, 24 and 25 show the epoch time and time to accuracy for GraphSAGE, GAT and GCN respectively for different sampling optimizations.
+Figure 13 shows the best accuracy achieved and 14 shows the time-to-accuracy plots for GraphSAGE, GAT and GCN models with different mini-batch optimizations (Neighbor sampler, ClusterGCN and SAINT sampler).
+Table 18. Best accuracies (%) for GAT model
+Pubmed Ogbn-arxiv Reddit Ogbn-products
+FG $77.70\pm 0.50$ $71.74\pm 0.22$ $95.17\pm 0.20$ $74.43\pm 0.05$
+PipeGCN $77.60\pm 0.37$ $70.34\pm 0.10$ $93.63\pm 0.07$ $74.08\pm 0.12$
+BNS-GCN $77.70\pm 0.10$ $71.75\pm 0.70$ $94.23\pm 0.02$ $74.50\pm 0.26$
+NS $78.50\pm 0.03$ $70.63\pm 0.12$ $95.58\pm 0.13$ $74.57\pm 0.07$
+ClusterGCN $80.07\pm 0.19$ $63.30\pm 0.09$ $92.49\pm 0.75$ $73.27\pm 0.29$
+GraphSaint $82.76\pm 0.95$ $72.58\pm 0.17$ $96.96\pm 0.06$ $70.47\pm 0.16$
+Table 19. Best accuracies (%) for GCN model
+Pubmed Ogbn-arxiv Reddit Ogbn-products
+FG $79.40\pm 0.12$ $71.54\pm 0.21$ $94.67\pm 0.20$ $75.71\pm 0.18$
+PipeGCN $77.00\pm 0.81$ $69.64\pm 0.27$ $93.91\pm 0.10$ $74.25\pm 0.34$
+BNS-GCN $77.70\pm 0.71$ $70.75\pm 0.36$ $93.23\pm 0.09$ $74.50\pm 0.32$
+NS $78.70\pm 0.23$ $71.00\pm 0.16$ $94.41\pm 0.14$ $79.31\pm 0.04$
+Cluster $77.50\pm 1.98$ $63.24\pm 0.10$ $92.42\pm 0.03$ $73.71\pm 0.08$
+GraphSaint $81.70\pm 0.71$ $69.72\pm 0.08$ $96.88\pm 0.06$ $72.65\pm 0.13$
 Table 20.
@@
 System # of GPUs ET TTA TTA Speed-up ET TTA TTA Speed-up ET TTA TTA Speed-up ET
-PipeGCN (Sync) 1 2.225 979.09 1.0x 5.857 5622.53 1.0x 3.510 1404.12 1.0x OOM
+Full-Graph 1 2.225 979.09 1.0x 5.857 5622.53 1.0x 3.510 1404.12 1.0x OOM
 2 1.514 665.98 1.5x 4.035 3873.98 1.5x 3.381 1352.32 1.0x 2.726
@@
 4 1.072 471.55 2.1x 2.890 2774.78 2.0x 2.452 980.60 1.4x 2.323
-PipeGCN (Async) 1 2.218 998.01 1.0x 5.856 5856.10 1.0x 3.503 1331.10 1.0x OOM
+PipeGCN 1 2.218 998.01 1.0x 5.856 5856.10 1.0x 3.503 1331.10 1.0x OOM
 2 1.334 600.08 1.7x 3.269 3269.20 1.8x 2.762 1049.37 1.3x 2.479
@@
 4 0.805 362.03 2.8x 1.879 1879.10 3.1x 1.523 578.78 2.3x 1.499
-BNS-GCN (Sync) 1 2.337 841.43 1.0x 6.232 6231.60 1.0x 4.010 842.10 1.0x OOM
+BNS-GCN 1 2.337 841.43 1.0x 6.232 6231.60 1.0x 4.010 842.10 1.0x OOM
 2 1.345 484.02 1.7x 3.472 3472.20 1.8x 3.466 727.82 1.2x 2.860
@@
 4 0.778 280.08 3.0x 2.193 2192.70 2.8x 1.710 359.06 2.3x 1.583
-DGL (Sync) 1 15.401 154.01 1.0x 17.468 262.01 1.0x 18.394 367.88 1.0x 22.465
+DGL 1 15.401 154.01 1.0x 17.468 262.01 1.0x 18.394 367.88 1.0x 22.465
 2 8.888 88.88 1.7x 8.575 128.62 2.0x 9.328 186.57 2.0x 11.580
@@
 4 4.627 46.27 3.3x 4.594 68.91 3.8x 4.715 99.02 3.7x 6.444
-Quiver (Sync) 1 16.230 162.30 1.0x 18.914 283.71 1.0x 17.457 349.15 1.0x 24.222
+Quiver 1 16.230 162.30 1.0x 18.914 283.71 1.0x 17.457 349.15 1.0x 24.222
 2 8.286 91.15 1.8x 9.002 135.03 2.1x 8.783 175.66 2.0x 14.134
@@
 4 0.260 2.60 3.5x 0.731 3.66 3.4x 0.643 9.65 3.8x 1.743
+Table 22. Time to accuracy (TTA) and epoch time (ET) for GraphSage and ogbn-papers100M (3 PCIe hosts)
+System Time (s)
+Epoch time (s) Full-Graph 22.00
+PipeGCN 1.03
+BNS-GCN 2.46
+DistDGL 38.37
+Time to accuracy (s) Full-Graph 19141.48
+PipeGCN 956.34
+BNS-GCN 859.26
+DistDGL 767.38
+Appendix A Ease of Hyperparameter Search
+Figure 10. Sensitivity to the number of layers and hidden size for GraphSAGE model.
+Figure 11.
+Ablation study with varying number of layers and hidden size for GCN model.
+Legend represents the size of each hidden layer.
+The X-axis shows the number of layers.
+The Y-axis shows the test accuracy.
+The top row is for mini-batch training and the bottom row is for full graph training.
+We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
+Figure 12.
+Ablation study with the varying number of layers and hidden size for GAT model.
+Legend represents the size of each hidden layer.
+The X-axis shows the number of layers.
+The Y-axis shows the test accuracy.
+The top row is for mini-batch training and the bottom row is for full graph training.
+We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
+How sensitive are the two training methods to changes in the hyperparameters?
+We evaluate the ease of finding a good hyperparameter setting using the two training approaches by measuring how accuracy varies as we vary the model architecture hyperparameters, i.e., the number of hidden layers and the size of the hidden vertex dimensions.
+We trained each model for a maximum of 1000 epochs, stopping earlier if the validation accuracy doesn’t increase for 50 epochs.
+Figure 10 illustrates the best test accuracy when training the GraphSAGE model, the top row shows mini-batch training experiments and the bottom row shows synchronous full-graph training.
+Notably, mini-batch training demonstrates remarkable consistency across varying model architectures.
+In contrast, full-graph training exhibits a wider range of variations in test accuracy when the architecture changes.
+We observed similar results for GCN and GAT models, which are reported in Figure 11 and Figure 12.
+Figure 10 and the other results for GCN and GAT, clearly show the effect of over-smoothing, which appears in GNNs at increased depths (Rusch et al., 2023).
+Mini-batch training on GraphSAGE does not show smoothing with up to 8 layers in Figure 10.
+In a set of separate experiments, we found that over-smoothing does occur with mini-batch training, but only after we increase the number of layers to $16$ or higher.
+Our experiments for GCN and GAT show that over-smoothing occurs at lower depths for mini-batch training but the resulting accuracy drop is still much less than that with full-graph training (see the Appendix).
+Takeaway.
+These results highlight the robustness of mini-batch training to alterations in model architecture.
+Full-graph training appears to be more sensitive to changes in architecture, so it requires a more careful and potentially expensive hyperparameter search.
+Whereas mini-batch training is consistently less sensitive to changes of those parameters than full-graph training, across all the GNN models we consider.
+In addition, mini-batch training is much less prone to the over-smoothing problem (Rusch et al., 2023) that affects deep GNN models with many layers.
+Table 23. Time to accuracy and epoch time for GraphSAGE
+Sampler pubmed ogbn-arxiv reddit ogbn-products
+Epoch time (s) Cluster GCN 0.0118 0.0254 0.2798 0.1212
+NS 0.0120 0.1592 2.6728 0.5552
+Saint Sampler 0.0123 0.1029 0.4689 0.0908
+Time to accuracy (s) Cluster GCN 2.5418 9.0578 97.6485 111.46
+NS 1.1168 48.3875 192.440 122.71
+Saint Sampler 1.0939 61.2358 151.910 42.9602
+Table 24. Time to accuracy and epoch time for GAT
+Sampler pubmed ogbn-arxiv reddit ogbn-products
+Epoch time (s) Cluster GCN 0.0353 0.0621 1.3446 0.4893
+NS 0.0222 0.2927 18.5336 1.2580
+Saint Sampler 0.0224 0.1176 0.3653 0.0524
+Time to accuracy (s) Cluster GCN 2.3293 15.3941 599.686 1109.32
+NS 3.8463 46.2388 1853.36 159.769
+Saint Sampler 2.4866 47.759 115.419 20.4534
+Table 25. Time to accuracy and epoch time for GCN
+Sampler pubmed ogbn-arxiv reddit ogbn-products
+Epoch time (s) Cluster GCN 0.0112 0.0259 0.2333 0.1030
+NS 0.0133 0.1814 4.4802 0.6186
+Saint Sampler 0.0126 0.1016 0.4506 0.0905
+Time to accuracy (s) Cluster GCN 2.5933 9.1563 39.6631 47.0832
+NS 1.8819 27.2038 488.345 64.3342
+Saint Sampler 2.5587 51.6232 159.962 29.0637
+Figure 13. Test accuracy achieved for the best architecture for GraphSAGE, GAT and GCN models with NeighborSampler, ClusterGCN and SAINT Sampler.
+Figure 14. Time to accuracy plots for GraphSAGE, GAT and GCN models with NeighborSampler, ClusterGCN and SAINT Sampler.

### v4

@@
-
+Graph Neural Network Training Systems: A Performance Comparison of Full-Graph and Mini-Batch
 Abstract.
@@
 We find that the mini-batch training systems consistently converge faster than the full-graph training ones across multiple datasets, GNN models, and system configurations.
-We also find that mini-batch training techniques converge to similar or often higher accuracy values as full-graph training ones, showing that mini-batch sampling is not necessarily detrimental to accuracy.
+We also find that mini-batch training techniques converge to similar to or often higher accuracy values than full-graph training ones, showing that mini-batch sampling is not necessarily detrimental to accuracy.
 Our work highlights the importance of comparing systems across different classes, using time-to-accuracy rather than epoch time for performance comparison, and selecting appropriate hyperparameters for each training method separately.
@@
 Samples are typically large, so sampling them and loading them into GPUs is expensive.
-The common rationale for using full-graph GNN training is to ensure that the training algorithm sees all the required dependent data in the graph, which is assumed to be the key to achieving high accuracy (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019).
+The common rationale for using full-graph GNN training is to ensure that the training algorithm sees all the required dependent data in the graph, which is assumed to be the key to achieving high accuracy (Thorpe et al., 2021; Wan et al., 2022b, a, 2023c; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023d; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019).
 Mini-batch GNN training introduces additional variability in the gradient estimation, even compared to regular mini-batch DNN training.
@@
 In this paper, we address this open issue by performing a thorough performance comparison of representative state-of-the-art GNN training systems across the two classes.
-For full-graph training, we consider PipeGCN (Wan et al., 2022b), BNS-GCN (Wan et al., 2022a), and AdaQP (Wan et al., [n.d.]b).
+For full-graph training, we consider PipeGCN (Wan et al., 2022b), BNS-GCN (Wan et al., 2022a), and AdaQP (Wan et al., 2024).
 For mini-batch training, we consider DGL (Wang et al., 2019), DistDGL (Zheng et al., 2020), and Quiver (Tan et al., 2023), and three sampling algorithms: Neighborhood Sampling (Hamilton et al., 2017), ClusterGCN (Chiang et al., 2019), and GraphSaint (Zeng et al., 2019).
@@
 These results are consistent with the evidence in the literature.
-Several papers that propose full-graph training systems use epoch time, which is the average time needed to execute one epoch, as the main metric to compare with mini-batch training systems, showing that the latter has a larger epoch time  (Wang et al., 2022; Wan et al., 2023b; Md et al., 2021; Chen et al., 2023).
+Several papers that propose full-graph training systems use epoch time, which is the average time needed to execute one epoch, as the main metric to compare with mini-batch training systems, and show that mini-batch systems have a larger epoch time  (Wang et al., 2022; Wan et al., 2023c; Md et al., 2021; Chen et al., 2023).
 Our evaluation confirms these results.
@@
 From a methodological perspective, our work highlights the importance of comparing GNN training systems across different classes and it also indicates a set of principles on how to perform this comparison.
+The source code, data, and other artifacts used for this evaluation are publicly available (Bajaj et al., 2024).
 2. Background and Motivation
@@
 Table 1. Review of literature reporting claims of how mini-batch and full-graph training compare. (Notation: TTA - time to accuracy, ACC - accuracy, FG - Full graph, MB - mini-batch, ET - Epoch time)
-Proposes System/paper Main FG vs. MB Claims Experimental evidence Year Open Source Evaluation Setup
+Proposes System / Paper Main FG vs. MB Claims Experimental Evidence Year Evaluation Setup
 Full-graph GNNPipe (Chen et al., 2023) FG lower ET ET 2023 Multi-host multi-GPU
@@
 ADGNN (Song et al., 2023a) FG higher ACC, FG better convergence 2023 Multi-host CPU, Multi-host single-GPU
-G3 (Wan et al., 2023b) FG higher ACC, lower ET ACC, ET 2023 Multi-host multi-GPU
-AdaQP (Wan et al., 2023c) MB information loss 2023 Multi-host multi-GPU
+G3 (Wan et al., 2023c) FG higher ACC, lower ET ACC, ET 2023 Multi-host multi-GPU
+AdaQP (Wan et al., 2023d) MB information loss 2023 Multi-host multi-GPU
 PipeGCN (Wan et al., 2022b) FG higher ACC ACC 2022 Multi-host multi-GPU
@@
 In this paper, we focus our evaluation on GPU-based training on multi-GPU systems, which is a common configuration for scalable GNN training.
-For a comprehensive and detailed discussion on distributed GNN training systems, see also (Shao et al., 2024; Lin et al., 2023).
+We consider system-level optimizations of the vanilla full-graph and mini-batch training pipelines that improve performance and support any GNN model, as defined in Section 2.1.
+We show the classes of optimizations we consider and the related baselines in Figure 2.
+For a comprehensive discussion on distributed GNN training systems, see also (Shao et al., 2024; Lin et al., 2023).
 Full-graph training systems.
@@
 The main focus of work on full-graph training systems has been on reducing the cost of vertex feature communication.
-We now discuss existing full-graph GNN training systems work in terms of two design dimensions: training data management and algorithmic optimizations (see Figure 2).
-Some full-graph training systems proposed optimizations that focused on the management of training data and do not introduce approximations that can potentially impact accuracy.
+We now discuss existing work on full-graph GNN training systems in terms of two design dimensions: training data management and algorithmic optimizations (see Figure 2).
+Some full-graph training systems focused on optimizing the management of training data without introducing approximations that can potentially impact accuracy.
 If there are not enough GPUs to store each partition in a different GPU, it is necessary to transfer data between the host memory and GPU memory.
-Prior work such as NeuGraph (Ma et al., 2019), RoC (Jia et al., 2020), G3 (Wan et al., 2023b), and HongTu (Wang et al., 2023) have proposed techniques to optimize host-GPU communication.
-When there are enough GPUs to distribute all the partitions of the graph in GPU memory, training proceeds as depicted in Figure 1(a) (Wang et al., 2022; Wan et al., 2023b; Cai et al., 2021; Tripathy et al., 2020).
+Prior works such as NeuGraph (Ma et al., 2019), RoC (Jia et al., 2020), G3 (Wan et al., 2023c), and HongTu (Wang et al., 2023) have proposed techniques to optimize host-GPU communication.
+When there are enough GPUs to distribute all the partitions of the graph in GPU memory, these optimizations are not required and training proceeds as depicted in Figure 1(a) (Wang et al., 2022; Wan et al., 2023c; Cai et al., 2021; Tripathy et al., 2020).
 Some recent works improve performance by further introducing algorithmic optimizations that can impact accuracy, unlike the work we described previously.
@@
 ADGNN proposes an aggregation-difference sampling algorithm (Song et al., 2023a).
-Finally, the AdaQP system proposed message quantization to reduce the communication cost, together with new partitioning algorithms (Wan et al., 2023c).
+Finally, the AdaQP system proposed message quantization to reduce the communication cost, together with new partitioning algorithms (Wan et al., 2023d).
 Figure 2. Classes of optimizations that impact performance (P) and accuracy (A), and representative systems evaluated in this work.
@@
 A significant amount of work on GNN training do not compare or evaluate their systems against systems using a different training approach and training pipeline (full-graph or mini-batch) (Kaler et al., 2023; Sun et al., 2023; Cai et al., 2023; Chen et al., 2018a; Hamilton et al., 2017; Chen et al., 2018b; Huang et al., 2018; Zeng et al., 2019; Gasteiger et al., 2022; Wang et al., 2021; Zhu et al., 2019; Zhang et al., 2020; Kipf and Welling, 2017; Gandhi and Iyer, 2021).
-Some prior work focuses only on the data-parallel mini-batching pipeline of Figure 1(b) and compares the performance of different low-level data management optimizations on top of the same underlying system (Yuan et al., 2024).
+Some prior work focused on the mini-batch training pipeline of Figure 1(b), evaluating the impact of using different data partitioning algorithms, configurations of the neighborhood sampling algorithm, and data loading optimizations within the same system (Yuan et al., 2024).
 In the following, we review 33 papers on GNN training that make claims comparing systems in the two classes.
@@
 Conflicting claims about how the two approaches compare.
-Several papers proposing full-graph systems (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wang et al., 2023; Song et al., 2023a) demonstrate that their systems achieve higher accuracy and lower epoch time compared to mini-batch systems.
-Some papers mention that mini-batch causes information loss, noisy gradients, and does not guarantee convergence (Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019; Wang et al., 2023; Song et al., 2023a) However, papers proposing mini-batch systems (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020) and some other works (Hu et al., 2020) present contradicting evidence and claims showing that mini-batch training leads to higher accuracy and faster convergence to the target accuracy level.
+Several papers proposing full-graph systems (Thorpe et al., 2021; Wan et al., 2022b, a, 2023c; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wang et al., 2023; Song et al., 2023a) demonstrate that their systems achieve higher accuracy and lower epoch time compared to mini-batch systems.
+Some papers mention that mini-batch causes information loss, noisy gradients, and does not guarantee convergence (Wan et al., 2023d; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019; Wang et al., 2023; Song et al., 2023a) However, papers proposing mini-batch systems (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020) and some other works (Hu et al., 2020) present contradicting evidence and claims showing that mini-batch training leads to higher accuracy and faster convergence to the target accuracy level.
 Other works claim full-graph training is impractical, requires high memory, and is hard to scale (Ai et al., 2023; Yang et al., 2022b; Zheng et al., 2022a; Chiang et al., 2019; Yao and Li, 2021).
@@
 Comparing performance only in terms of epoch time.
-Certain works compare full-graph and mini-batch training systems based only on epoch time (Wang et al., 2022; Wan et al., 2023b; Md et al., 2021; Chen et al., 2023).
+Certain works compare full-graph and mini-batch training systems based only on epoch time (Wang et al., 2022; Wan et al., 2023c; Md et al., 2021; Chen et al., 2023).
 As we will show, a comparison in terms of time-to-accuracy is more informative given the differences in the two training approaches.
@@
 This conflates the effect of the training approach with the GNN model choice, since different models may exhibit varying performance even when trained with the same technique.
-Some papers do not mention explicitly whether they perform separate hyperparameter tuning for each method (Wan et al., 2023b, 2022b, 2022a; Thorpe et al., 2021; Zhao et al., 2021; Kaler et al., 2022).
+Some papers do not mention explicitly whether they perform separate hyperparameter tuning for each method (Wan et al., 2023c, 2022b, 2022a; Thorpe et al., 2021; Zhao et al., 2021; Kaler et al., 2022).
 As we show in our experimental evaluation of Section 5, given a dataset and GNN model, it is possible to find hyperparameter settings where one method achieves higher accuracy than another and vice versa.
@@
 All systems adopt a pipeline similar to the one of Figure 1(a).
-The Full-Graph baseline is synchronous and does not use any optimization that can impact accuracy (Wan et al., 2023a).
-In the dimension of algorithmic optimization, we consider three baselines representing three classes of algorithmic optimizations: PipeGCN for asynchronous training (Wan et al., 2022b, 2023a), BNS-GCN for sampling (Wan et al., 2022a, [n.d.]a), and AdaQP for message quantization (Wan et al., 2023c, [n.d.]b).
+The Full-Graph baseline is synchronous and does not use any optimization that can impact accuracy (Wan et al., 2023b).
+In the dimension of algorithmic optimization, we consider three baselines representing three classes of algorithmic optimizations: PipeGCN for asynchronous training (Wan et al., 2022b, 2023b), BNS-GCN for sampling (Wan et al., 2022a, 2023a), and AdaQP for message quantization (Wan et al., 2023d, 2024).
 For mini-batch training, we consider three baselines: DGL (Wang et al., 2019; dgl, 2023), its distributed version DistDGL (Zheng et al., 2020; dgl, 2023), and Quiver (Tan et al., 2023; Quiver, 2023).
@@
 DistDGL partitions the training dataset across multiple hosts.
-In terms of GPU caching, DGL only supports local GPU caching while Quiver also supports distributed GPU caching.
-Quiver uses distributed input feature caching for the Orkut and Papers100M datasets, which are too large for local caching.
+All systems use GPU caching whenever possible.
+DGL only supports local GPU caching while Quiver also supports distributed GPU caching, which is necessary to cache the Orkut and Papers100M datasets.
 Mini-batch training can use different sampling algorithms to select a subset of the k-hop neighbors of the vertices in the mini-batch.
@@
 Different mini-batch sampling algorithms.
-Finally, we combine DGL with different mini-batch sampling algorithms than Neighborhood Sampling (NS).ClusterGCN achieves on average 2.8×2.8$\times$2.8 × faster TTA than NS across datasets, ranging from 0.14×0.14$\times$0.14 × slower to 12.3×12.3$\times$12.3 × faster depending on the dataset, and its ET is 6.4×6.4$\times$6.4 × faster on average.
+Finally, we combine DGL with different mini-batch sampling algorithms than Neighborhood Sampling (NS).
+ClusterGCN achieves on average 2.8×2.8$\times$2.8 × faster TTA than NS across datasets, ranging from 0.14×0.14$\times$0.14 × slower to 12.3×12.3$\times$12.3 × faster depending on the dataset, and its ET is 6.4×6.4$\times$6.4 × faster on average.
 GraphSaint has 3.2×3.2$\times$3.2 × faster TTA than NS on average, ranging from 0.5×0.5$\times$0.5 × slower to 16.1×16.1$\times$16.1 ×, and faster ET by 9.3×9.3$\times$9.3 × on average (see appendix tables 24 and 25).
@@
 For this experiment, we replace DGL with DistDGL, its distributed version.
-We report the hyperparameters in the appendix in table 15.
+We report the hyperparameters in the in the appendix table 15.
 We consider the GraphSAGE model because it is the only one implemented in the DistDGL distribution.
@@
 Even though a system takes longer than another to reach the same target test accuracy, it could still be preferable to use it if it can converge to a higher test accuracy.
-Indeed, some works on GNN systems assert that full-graph training achieves higher accuracy and that mini-batch training does not guarantee convergence (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
+Indeed, some works on GNN systems assert that full-graph training achieves higher accuracy and that mini-batch training does not guarantee convergence (Thorpe et al., 2021; Wan et al., 2022b, a, 2023c; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023d; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
 Figure 5 summarizes the test accuracy values reported in the literature we reviewed for different GNN models and datasets.
@@
 The figure illustrates a wide distribution of accuracy values for the same dataset and model combination.
-In this Section, we shed light onto this apparent contradiction and answer two questions: Q3) What is the test best accuracy we can achieve using full-graph and mini-batch training?
+In this Section, we shed light onto this apparent contradiction and answer two questions: Q3) What is the best test accuracy we can achieve using full-graph and mini-batch training?
 Q4) What test accuracy can we achieve when using a training method with the best hyperparameters found for the other method?
-Different training systems have different optimizations, so Q5: What is the impact of the optimizations on accuracy?
+Different training systems have different optimizations (see Figure 2), so Q5: What is the impact of the optimizations on accuracy?
 Figure 5.
@@
 Full-graph training has lower accuracy than mini-batch training and the gap between the two methods is larger, reaching $6.22\%$ for the reddit dataset.
-Our evaluation of GAT and GCN is reported on the appendix in table 9 and table 10 which leads to similar conclusions.
+Our evaluation of GAT and GCN is reported in the appendix in table 9 and table 10 which leads to similar conclusions.
 For GAT, the largest difference between FG-train and MB-train is $2.85\%$ in favor of MB-search for arxiv and $1.9\%$ in favor of FG-search for pubmed.
@@
 Impact of optimizations on accuracy.
-Many performance optimizations we consider can result in a different accuracy than the vanilla methods.
+Many performance optimizations we consider can result in a different accuracy than the vanilla methods (see Figure 2).
 This section answers the question Q5: What is the impact of the optimizations on accuracy?
@@
 The results are shown in Figure 9.
-The computational cost of sampling is noticeable but is still lower than the training cost.
+The computational cost of sampling is noticeable but is still lower than the training cost in most cases.
 It decreases when we increase the number of workers/partitions because each epoch requires fewer iterations.
 Takeaway.
-This analysis shows that vanilla full-graph training has a higher communication cost than vanilla mini-batch training, especially with larger graphs.
+This analysis shows that vanilla full-graph training has a higher communication cost than vanilla mini-batch training when the GNN model is not very deep.
 In practice, mini-batch training can replicate input features across multiple workers to eliminate or reduce communication costs.
 The high communication cost of full-graph training has been the focus of prior work on full-graph training systems, as discussed in Section 2.2.
-The analysis shows that using sampling in mini-batch training also results in a substantial reduction in computation cost compared to full-graph training systems, especially with larger graphs.
-Overcoming this gap while maintaining the full-graph training approach and avoiding sampling is challenging.
+The analysis shows that using sampling in mini-batch training also results in a substantial reduction in computation cost compared to full-graph training systems.
 Figure 9. Fraction of epoch time spent performing sampling in mini-batch training for GraphSage.
 7. Conclusion
-A common rationale for using full-graph training is to achieve high accuracy by avoiding the information loss results from sampling.
+Our work has evaluated various full-graph and mini-batch systems, encompassing a wide range of optimizations (see Figure 2).
+A common rationale for using full-graph instead of mini-batch training is to achieve higher accuracy by avoiding the information loss resulting from sampling.
 In our evaluation, we did not observe these expected benefits.
 Mini-batch training systems not only achieve a lower time-to-accuracy across all models, datasets, and hardware configurations we consider but also a comparable or even higher accuracy with proper hyperparameter tuning.
-Mini-batch training also has some limitations.
+Among the optimizations we considered for mini-batch training, GPU-based sampling and GPU caching should be used whenever feasible.
+The choice and configuration of the sampling algorithm is critical for performance and accuracy.
 The common Neighborhood Sampling algorithm is able to consistently achieve good accuracy, but it still shows a slightly larger variance between training runs than vanilla full-graph training.
 Other sampling algorithms we evaluated are not always able to converge to a high accuracy, depending on the dataset.
-Finding sampling algorithms that have consistently higher performance than Neighborhood Sampling is an interesting avenue of research.
-Our experiments show that as GNN models become deeper, the advantage of mini-batch training over full-graph training reduces.
 Distributed sampling over multiple hosts can also become a bottleneck, especially when the network is slow.
-An important advantage of mini-batch training systems is that the sampling algorithms are interchangeable components.
-This modularity and flexibility in choosing the right sampling algorithm is key since different algorithms perform very differently on different datasets.
-Our experiments show that providing efficient sampling implementations is important for multi-host setups, especially when using a slow network.
-In order to support future innovations in sampling algorithms, it is important to have systems that allow ML practitioners to implement new sampling algorithms and run them on GPUs, such as (Jangda et al., 2021; Pandey et al., 2020; Gong et al., 2023; Wang et al., 2021).
-In terms of improving full-graph training systems, our cost model analysis and single-GPU performance experiments reveal that in addition to the well-known communication cost of full-graph training—a bottleneck addressed by much previous work—there is a computation cost gap that is difficult to bridge without employing sampling.
-A natural open question is whether combining full-batch training with more aggressive sampling, forgoing full-graph aggregation, can be more effective than standard mini-batch training.
-In our hyperparameter tuning, however, we found that the accuracy of mini-batching typically peaks with batch sizes around 1024/4096 and does not grow further with larger values, so the advantages of using full-batch training are unclear.
+Systems that allow ML practitioners to implement new sampling algorithms and run them on GPUs, such as (Jangda et al., 2021; Pandey et al., 2020; Gong et al., 2023; Wang et al., 2021), can support future innovations in efficient sampling algorithms.
+For full-graph training systems, the algorithmic optimizations we considered can substantially improve performance with minimal or no impact on accuracy.
+Our performance evaluation and cost model analysis reveal that in addition to the well-known communication bottleneck addressed by much previous work, future optimizations should also reduce the computation cost gap with mini-batch training.
+A natural open question is whether combining the full-batch training pipeline of Figure 1(a) with more aggressive sampling, forgoing full-graph aggregation, can be more effective than mini-batch training.
+In our hyperparameter tuning, however, we found that the accuracy of mini-batching typically peaks with batch sizes around 1024/4096 and does not grow further with larger values, so the advantage of using full-batch training is unclear.
 Appendix